### v1

# Learning Continually by Spectral Regularization

Abstract
footnotetext: Department of Computing Science, University of Alberta
Department of Computer Science, Stanford University
Google DeepMind
Canada CIFAR AI Chair
Loss of plasticity is a phenomenon where neural networks become more difficult to train during the course of learning. Continual learning algorithms seek to mitigate this effect by sustaining good predictive performance while maintaining network trainability. We develop new techniques for improving continual learning by first reconsidering how initialization can ensure trainability during early phases of learning. From this perspective, we derive new regularization strategies for continual learning that ensure beneficial initialization properties are better maintained throughout training. In particular, we investigate two new regularization techniques for continual learning: (i) Wasserstein regularization toward the initial weight distribution, which is less restrictive than regularizing toward initial weights; and (ii) regularizing weight matrix singular values, which directly ensures gradient diversity is maintained throughout training. We present an experimental analysis that shows these alternative regularizers can improve continual learning performance across a range of supervised learning tasks and model architectures. The alternative regularizers prove to be less sensitive to hyperparameters while demonstrating better training in individual tasks, sustaining trainability as new tasks arrive, and achieving better generalization performance.
1 Introduction
A longstanding goal of machine learning research is to develop algorithms that can learn continually and cope with unforeseen changes in the data distribution [1, 2]. Current learning algorithms, however, struggle to learn from dynamically changing targets and are unable to adapt gracefully to unforeseen changes in the distribution during the learning process [3, 4, 5, 6]. Such limitations can be seen as a byproduct of assuming, one way or another, the problem is stationary. Recently, there has been growing recognition of the fact that there are limitations to what can be learned from a fixed and unchanging dataset [7], that there are implicit non-stationarities in many problems of interest [8], and that some real-world problems benefit from learning continually [9, 10].
The concept of plasticity has been receiving growing attention in the continual learning literature, where the loss of plasticity—a reduction in the trainability of the neural network [11, 5, 12]—has been noted as a critical shortcoming in current learning algorithms. Learning algorithms, and more specifically neural networks, that are performant in the non-continual learning setting often struggle when applied to continual learning problems. Settings where a neural network must continue to learn after changes occur in the observations exhibit a striking loss of plasticity such that learning slows down or even halts after successive changes in the learning distribution [13, 6, 14].
Several aspects of a learning algorithm have been found to contribute to, or mitigate, loss of plasticity. Examples include the type of optimizer [6], the step-size [15, 16], the number of optimiser iterations [5], and the use of specific regularizers [11, 14, 17]. Such factors hint that there might be simpler underlying optimisation principles that govern the loss of plasticity. For example, the success of several methods that regularize neural networks towards properties of the initialization suggests that some of those properties mitigate loss of plasticity [11, 17, 14].
Various properties of neural network weights at initialization have been shown to be associated with trainability. Previous analyses have demonstrated that the trainability of deep neural networks can be improved by ensuring that the variance of hidden activations and gradients remains uniform across the different layers of the neural network [18], which has the effect of keeping the average singular value of the layerwise Jacobians close to one. A stronger condition is dynamical isometry, where all singular values are close to one, or exactly one in the case of an orthogonal initialization [19, 20, 21]. Over the course of learning, the singular values deviate from their initialization and tend to grow over time [22], which can reduce trainability and impede continual learning.
Given the effectiveness of initialization in ensuring the trainability of a neural network, continual learning algorithms may benefit from sustaining the relevant properties of the weights. This is the motivation of several methods that use the initialization either through regularization [17] or weight reinitialization [11, 23, 6]. Such methods leverage initialization explicitly, either by regularizing parameters towards their initial values or by resetting parameters by resampling them from the initialization distribution. In this paper, we seek to more directly address loss of trainability by sustaining the key properties present at initialization, rather than replicate the initialization explicitly.
We investigate the key properties of the initialization that ensure trainability, how these properties are lost over the course of learning, and the effects this has on continual learning. Based on this analysis, we then introduce two regularizers that try to maintain the identified properties: we propose to use the Wasserstein distance to regularize the parameters towards the initialization distribution, which is less restrictive than regularizing towards initial weights, and we propose spectral regularization to control the maximum singular value of the weight matrix to directly address the deviation of the singular values from initialization over the course of learning. Our experiments show that spectral regularization is more performant and less sensitive to hyperparameters than other regularizers across datasets, nonstationarities, and architectures. While well-tuned regularizers are often able to mitigate loss of plasticity to a varying degree, learning continually with spectral regularization is robust, achieving high initial and sustained performance. Finally, we show that spectral regularization is also capable of improving generalization with both Vision Transformer [24] and ResNet-18 [25] on continual versions of SVHN2 [26], CIFAR10, and CIFAR100 [27].
2 Problem Setting
We investigate the trainability of learning algorithms, and specifically neural networks, in task-agnostic continual supervised learning. We assume the learning algorithm uses a variant of stochastic gradient descent to update a neural network with $L$ layers, defined recursively as $f_{\theta}(x):=h_{L}(x)$, where $h_{0}(x)=x$ is the input vector, $h_{l+1}(x)=\phi(\mathbf{W}_{l}h_{l}(x)+\mathbf{b}_{l})$ with an element-wise activation function $\phi$ and parameters $\theta=\{\theta_{L},\dots,\theta_{1}\}$ where $\theta_{l}=\{\mathbf{W}_{l},\mathbf{b}_{l}\}$ is the weight and bias parameter for layer $l$ [^1] In this notation we suppressed the presence of biases, which can effectively be included in the parameters. We assume that $f_{\theta}$ is trained over a sequence of tasks: the $\tau$th task is specified by a distribution $p_{\tau}$ over the observation-target pairs (denoted by $(x,y)$). For simplicity we assume that the tasks (data distribution) change periodically after every $T$ iterations, but we consider the task-agnostic setting, where the learning algorithm does not have access to information about the tasks except through the data that it samples. For each iteration in the task, $t\in\left((\tau-1)T,\tau T\right]$, the learning algorithm must optimise the neural network’s parameters to minimize the objective
$J_{\tau}(\theta)=\mathbb{E}_{(x,y)\sim p_{\tau}}\big{[}\ell(f_{\theta}(x),y)\big{]}$
for some loss function $\ell$. In this paper we consider gradient-based methods to learn the weight parameters; the basic version of gradient descent would update the parameters at iteration $t$ as $\theta^{(t+1)}=\theta^{(t)}-\alpha\nabla_{\theta}J_{\tau(t)}(\theta)\big{|}_{\theta=\theta^{(t)}}$, where $\tau(t)=\lfloor t/T\rfloor$ denotes the current task number. In practice this is done by considering the empirical error instead of $J_{\tau(t)}$ computed from samples $(x_{t},y_{t})\sim p_{\tau}$ for $t\in\left((\tau-1)T,\tau T\right]$ generated from the task distribution, and usually some more involved optimization algorithm.
The evaluation of the continual learning algorithm is performed at the end of a task [5]. Specifically, the learning algorithm is evaluated at times $t=\tau T$, after being given $T$ iterations to learn on task $\tau$. It is commonly assumed that each task is sampled independently and identically, and that each task is equally difficult, meaning that a neural network which is randomly initialized on each task is able to reach a similar objective value for any particular task [6, 12, 5].
Loss of plasticity in the continual learning literature can refer to either loss of trainability [11, 5] or to loss of generalization [15]. Because trainability is a requirement for learning and generalization, we focus on loss of trainability, using it interchageably with loss of plasticity. Specifically, we use loss of trainability to refer to the phenomenon that the objective value, $J_{\tau}(\theta^{(\tau T)})$, increases as a function of the task $\tau$. Equivalently, the performance measures, such as accuracy, decrease with new tasks. Performance can also decrease, then plateau, and in this case the network has lost trainability.
3 Spectral Properties and Continual Trainability
A growing body of work has identified that neural networks trained on continually changing distributions exhibit loss of trainability, slowing down (or even preventing) learning on new tasks [11, 13, 6, 5]. At a high level, the reason behind this phenomenon is simple: the solution of one task becomes the initialization for learning on the next task, and if this initialization is not sufficiently good to enable learning in the new task, we face the aforementioned problem. To avoid this issue, we need to keep the network parameters within a region that can serve as good initialization at the start of a task. This approach comes with two challenges: (i) determining a suitable region for parameter initialization; (ii) ensuring the initialization region does not prevent effective learning in the tasks encountered.
It is easy to satisfy (i) and (ii) in isolation: using a standard initialization algorithm, such as those of [28, 29], and then keeping the parameters fixed at this value satisfies (i) but clearly does not satisfy (ii). Conversely, not restricting the parameter space while using a standard learning algorithm addresses (ii) but not (i). We seek a balance of the two requirements. Note that (i) and (ii) are not binary properties, hence there can be a trade-off (e.g., allowing less favorable parameters for next task initialization allow better optimization in the current task). It is worth noting that such an approach has been successfully applied in the context of online convex optimization in changing environments, where several algorithms can be recognized as running a standard learning algorithm (such as mirror descent) on a carefully selected subset of the parameter space (see, e.g., [30, 31, 32]).
Our approach is the following: to address problem (i), we first identify key properties that initialization algorithms [28, 29] impose on the starting parameters of a neural network. Then, we propose two regularizers targeting these properties strikes a good balance between (i) and (ii).
3.1 Spectral properties at initialization
[…]
3.2 An illustrative example
To make these points more concrete, consider a single layer neural network $f_{\theta}(x)=\theta_{2}\texttt{ReLU}(\theta_{1}x)$ mapping $x\in\mathbb{R}^{2}$ to $\mathbb{R}$. Suppose that the first task is to fit $x=(1,0)^{\top}$ and $y=0$ in mean-squared error. An optimal solution with $f_{\theta}(x)=y$ is $\theta_{1}=\begin{bmatrix}1&0\\ 0&a\end{bmatrix}$ and $\theta_{2}=(0,a)$ for an arbitrarily small value of $a\geq 0$. Now consider that the next task is to fit $x=(0,1)^{\top}$ and $y=1$. The gradients of the loss $\ell(f_{\theta}(x),y)=\frac{1}{2}(y-\ell(f_{\theta}(x))^{2}$ are, for the given new $(x,y)$ pair, $\nabla_{\theta_{2}}\ell(f_{\theta}(x),y)=(f_{\theta}(x)-y)\texttt{ReLU}(\theta_{1}x)=(a^{2}-1)(0,a)^{\top}$ and $\nabla_{\theta_{1}}\ell(f_{\theta}(x),y)=(f_{\theta}(x)-y)x\texttt{ReLU}^{\prime}(\theta_{1}x)^{\top}=(a^{2}-1)\begin{bmatrix}0&0\\ 0&a\end{bmatrix}$. Performing updates with these gradients keeps the $\Theta(a)$ parameters at the same order while maintaining a loss of approximately 1. The condition number of the weight matrix $\theta_{1}$ is $\kappa=1/a$, which requires $\Theta(1/a)=\Theta(\kappa)$ update steps to make significant progress during learning. (A more detailed analysis of another example is given in Appendix A.2.)
3.3 Trainability and Effective Gradient Diversity
We can generalize the above observation to a problem of reduced gradient diversity. By gradient diversity, we mean the spread of singular vectors in the matrix of per-example stochastic gradients, $\mathbf{G}=[\mathbf{g}_{1},\ldots,\mathbf{g}_{m}]$. For this analysis, we focus on the rank, which is one particular summary statistic for the set of singular values that counts the number of non-zero singular values, $\text{rank}(\mathbf{G})=|\{i:\sigma_{i}(\mathbf{G})>0\}|$. However, for practical purposes the rank is problematic because it is unstable to perturbations [34, Theorem 1]. In our experiments below we will consider the condition number, $\nicefrac{{\sigma_{1}(\mathbf{G})}}{{\sigma_{m}(\mathbf{G})}}$ and effective rank, $\text{erank}(\mathbf{G})=\sum\bar{\sigma}_{i}(\mathbf{G})\log\bar{\sigma}_{i}(\mathbf{G})$ where $\bar{\sigma}_{i}(\mathbf{G})=\nicefrac{{\sigma_{i}(\mathbf{G})}}{{\sum_{i}\sigma_{i}(\mathbf{G})}}$ [35]. These measures are particularly sensitive to a growing largest singular value, if it grows faster than the smallest singular values (our experiments show this to be the case in Section 5.3). In this case, we say there is a reduction in the effective gradient diversity.
A reduction in the rank of the gradient matrix results in collinear gradients on different datapoints, limiting the diversity of gradient directions used in the updates. This can have an adverse effect on learning. In the extreme case, where the gradient matrix is rank one, every datapoint provides a gradient in the same direction, even if the datapoints correspond to different classes. Consider the gradient of the loss on a particular datapoint $(x_{i},y_{i})$ with respect to the weight matrix of a hidden layer, $\theta_{l}\in\mathbb{R}^{d\times d}$, which can be written recursively as $\mathbf{G}_{l,i}=\nabla_{\theta_{l}}\ell(f_{\theta}(x_{i}),y_{i})=\delta_{l,i}h_{l-1,i}^{\top}$, where $\delta_{l,i}=\theta_{l+1}^{\top}\delta_{l+1,i}\mathbf{D}_{l,i}$ is the error gradient from the previous layer with $\delta_{L,i}=\nicefrac{{\partial\ell(f_{\theta}(x_{i}),y_{i})}}{{\partial f_{\theta}}}$ and $\mathbf{D}_{l,i}=\text{Diag}(\phi^{\prime}([\theta_{l}h_{l,i}]_{1}),\dotso,\phi^{\prime}([\theta_{l}h_{l,i}]_{d}))$. We can rewrite the gradient in terms of its vectorization $\mathbf{g}_{l,i}=\text{vec}(\mathbf{G}_{l,i})=(\mathbf{I}_{d}\otimes\theta_{l+1}^{\top})\mathbf{v}_{l,i}$ where $\mathbf{I}_{d}$ is the $d\times d$ identity matrix and $\mathbf{v}_{1,i}=\text{vec}(\delta_{l+1,i}\mathbf{D}_{l,i}h_{l-1,i}^{\top})\in\mathbb{R}^{d^{2}}$ is the vectorization of data-dependent terms. The matrix of gradients for different datapoints is the concatenation of the per-example gradients, $\mathbf{G}_{l}=[\mathbf{g}_{l,1},\mathbf{g}_{l,2},\dotso,\mathbf{g}_{l,m}]=(\mathbf{I}_{d}\otimes\theta_{l+1}^{\top})\mathbf{V}_{l}$ where $\mathbf{V}_{l}=[\mathbf{v}_{l,1},\mathbf{v}_{l,2},\dotso,\mathbf{v}_{l,m}]$. The rank of $\mathbf{G}_{l}$ is upper bounded as
$\text{rank}(\mathbf{G}_{l})<\min\{\text{rank}(\mathbf{I}_{d}\otimes\theta_{l+1}^{\top}),\text{rank}(\mathbf{V}_{l})\}=\min\{d\,\text{rank}(\theta_{l+1}^{\top}),\text{rank}(\mathbf{V}_{l})\}.$
Thus, if the (effective) rank of $\theta_{l+1}$ decreases, then the (effective) rank of the gradient matrix may decrease as well. The rank of the gradient matrix can even decrease due to rank reduction in parameters at other layers, or through the rank of the representation, which others have noted to occasionally correlate with loss of trainability [6, 5, 17].
3.4 Spectral Properties Away From Initialization During Continual Learning
It has been observed empirically that the parameter norm grows at a $\sqrt{t}$ rate in large-scale self-supervised learning [36]. Similar observations have been made in continual learning, provided that the neural network does not stop learning due to loss of trainability [13, 6, 37]. A growing parameter norm is problematic from an optimisation perspective. Specifically, the parameter norm under consideration is the Frobenius norm, equal to the sum of squared singular values, $\|\mathbf{\theta}\|_{F}^{2}=\sum_{i}\sum_{j}[\theta]_{ij}^{2}=\sum_{i}\sigma_{i}(\theta)^{2}$. Thus, a growing parameter Frobenius norm is equivalent to an increasing sum of the squared singular values. In particular, since $\|\theta_{l}\|_{F}\leq\sqrt{\text{rank}(\theta_{l})}\sigma_{max}(\theta_{l})$ where $\sigma_{max}(\theta_{l})$ denotes the spectral norm (defined as the largest singular value) [38], if the parameter (Frobenius) norm of layer $l$ grows at a $\sqrt{t}$ rate, then the spectral norm of the parameter matrix $\theta_{l}$ also increases at the same rate. This leads to an increase in the spectral norm of the layerwise Jacobian, $\sigma_{d}(\mathbf{D}_{l})\sigma_{max}(\theta_{l})\leq\sigma_{max}(\mathbf{J}_{l})$, which may harm trainability. Moreover, as we argued in the previous section, an increase in the spectral norm can result in reduced effective gradient diversity.
4 Regularizing for Trainability in Continual Learning
If important properties of the initialization are lost during the course of learning, it is natural to regularize the neural network towards the initialization. This is the motivation for regularization [17] and weight reinitialization [23, 6] as loss of plasticity mitigators. However, our motivation is to more directly target key properties of initialization, using the insights from Section 3.
To directly preserve key initialization properties, we will develop regularizers $\mathcal{R}_{\tau(t)}(\theta,s)$ that are a function of (i) the parameters of the neural network, $\theta$, (ii) the data, through the current task $\tau(t)$, and of (iii) auxiliary information, such as the parameters at initialization, through the state variable $s$. Naturally, not every regularizer takes all of these elements into account. The regularizer is optimized alongside the base objective, $J_{\tau(t)}(\theta)$. We write the composite objective as $J^{\lambda}_{\tau(t)}(\theta)=J_{\tau(t)}(\theta)+\lambda\mathcal{R}_{\tau(t)}(\theta,s)$, where $\lambda$ is a tunable hyperparameter governing the regularization strength.
One commonly used regularizer in both continual and non-continual learning is L2 regularization, $\mathcal{R}_{\tau(t)}(\theta,\varnothing)=\|\theta\|^{2}$, which regularizes the parameters towards zero. A recent alternative, L2 regularization towards the initialization, $\mathcal{R}_{\tau(t)}(\theta,\theta^{(0)})=\|\theta-\theta^{(0)}\|^{2}$, was proposed to deal with sensitivity to parameters near zero [17]. Regularizing towards the particular parameters present at initialization allows the neural network to regenerate parameters, providing a soft reset to parameters if the gradient on the base objective is zero (see Appendix A.3 for an example and Appendix A.5 for more details). One potential problem with L2 regularization towards the initialization is that it may prevent the parameters from deviating from the particular sampled value from the initialization distribution.
4.1 Wasserstein Regularization Towards Initialization
A first attempt to preserve the properties of the initialization is to regularize the parameter distribution rather than their specific values. Thus, we first propose a regularization that is simple and penalizes the parameters if they are far from the initialization distribution. At initialization, the parameters at layer $l$ are sampled i.i.d. $\theta_{i,j}\sim p^{(l,0)}(\theta)$ according to some pre-determined distribution $p^{(l,0)}$, such as the Glorot initialization [18] The distribution of parameters is only known at the initialization (at iteration 0); for any particular layer, the distribution $p^{(l,t)}$ is no longer has a known closed form at iteration $t$ during training. However, it is still possible to regularize the empirical distribution towards the initialization distribution by using the empirical Wasserstein-2 distance, defined as the difference in the order statistics [39]. We denote the flattened parameter matrix for layer $l$ at iteration $t$ by $\bar{\theta}^{(l,t)}$, and its order statistics by $\theta_{(i)}^{(l,t)}$ for $i=1,\ldots,d_{l}$, representing the $i$-th smallest parameter at iteration $t$ for layer $l$ (here $d_{l}$ denotes the number of parameters at layer $l$). The Wasserstein regularizer uses the empirical Wasserstein-2 distance for each layer of the neural network:
$\mathcal{R}_{\tau(t)}(\theta,\theta^{(0)})=\sum_{l=1}^{L}\sum_{i=1}^{d_{l}}\left(\bar{\theta}_{(i)}^{(l,t)}-\bar{\theta}_{(i)}^{(l,0)}\right)^{2}.$
In the above equation, we are taking the L2 difference between the order statistics of each layer’s parameters at initialization and at iteration $t$ during training. One of the benefits of using this distributional distance is that the parameters have more leeway to move away from the initialization compared to the L2 regularizer (see Appendix A.6 for details).
4.2 Spectral Regularization for Continual Learning
In Section 3, we argued that the optimisation dynamics of continual learning may impede the continued trainability of a neural network. In particular, we showed that a growing parameter norm results in a growing spectral norm and condition number which can harm trainability by reducing effective gradient diversity. Now we investigate regularization as a means of controlling the spectral norm to maintain trainability of the neural network.
The parameter matrix $\theta_{l}$ in layer $l$ is a composition of the weight matrix $\mathbf{W}_{l}$ and the bias $\mathbf{b}_{l}$, giving the augmented parameter matrix $\theta_{l}=[\mathbf{W}_{l}|\mathbf{b}_{l}]$ (with this the notation the equation defining the of the neural network $f_{\theta}$ can be rewritten as $h_{l+1}(x)=\phi(\mathbf{W}_{l}h_{l}(x)+\mathbf{b}_{l})$). As discussed in Section 3, our goal is to control the maximum singular value of $\theta_{l}$. Using that $\sigma_{max}(\theta_{l})=\sigma_{max}([\mathbf{W}_{l}|\mathbf{b}_{l}])\leq\sigma_{max}(\mathbf{W}_{l})+\sigma_{max}(\mathbf{b}_{l})$, we achieve this by regularizing the spectral norm of each parameter in the layer separately: the spectral norm of the multiplicative weight parameter $\mathbf{W}_{l}$ is regularized towards 1 and the spectral norm of the additive bias parameter $\mathbf{b}_{l}$ is regularized towards 0:
$\mathcal{R}_{\tau(t)}(\theta_{l})=\sum_{l\in\text{layers}}\left((\sigma_{max}(\mathbf{W}_{l})^{k}-1)^{2}+\left(\sigma_{max}(\mathbf{b}_{l})^{k}\right)^{2}\right)=\sum_{l\in\text{layers}}\left((\sigma_{max}(\mathbf{W}_{l})^{k}-1)^{2}+\|\mathbf{b}_{l}\|_{2}^{2k}\right).$
We add the exponent $k$ to the spectral norm to penalize large spectral norms deviating from one. In our experiments, we set $k=2$ and provide an ablation study in Appendix C.2. The maximum singular value can be computed efficiently using power iteration [40, 41]. Similarly to previous work using spectral regularization for generalization in supervised learning, we find that a single iteration is sufficient for effective regularization [42].
Spectral regularization explicitly regularizes the spectral norm (the maximum singular value) of each layer and addresses the problem of reduced effective gradient diversity described in Section 3.3. We only minimize the maximum singular value. We found that the smallest singular value remains relatively constant across all the architectures that we consider (see Appendix C.5). For implementation details for other layers, such as normalization layers and convolutional layers, see Appendix A.9.
In addition to the exponent $k$, our approach to spectral regularization is different in two important ways compared to spectral-norm regularization [42]. First, we regularize the spectral norm of every parameter, including bias terms and normalization parameters (see Appendix A.9). This is required because every parameter experiences norm growth. Second, we regularize the spectral norm of the multiplicative parameters to one, rather than zero. Aggressive regularization of the spectral norm towards zero can lead to collapse issues similar to L2 regularization towards zero. These two differences are crucial to improving trainability in continual learning. Lastly, spectral regularization is preferable over spectral normalization for two reasons: (i) spectral normalization is data-dependent (see Section 2.1 and Equation 12 in [43], and Appendix A.8), which can be problematic in continual learning due to the changing data distribution, and (ii) other forms of normalization are already used to train deep neural networks, such as LayerNorm [44], and it has been shown that spectral normalization does not improve trainability [5]. See Appendix A.7 for more details.
5 Experiments
Our experiments are designed to investigate the effect of both Wasserstein and spectral regularization on trainability in continual supervised learning. They cover several different datasets, types of nonstationarity, and architectures. We compare our proposed regularizers against baselines that have shown to improve trainability in previous work, which we detail below. Overall, our experiments demonstrate that spectral regularization, and to a lesser extent Wasserstein regularization (i) consistently mitigates loss of trainability on a wide variety of continual supervised learning problems, including training large neural networks for thousands of epochs across hundreds of tasks; (ii) is highly robust to the number of training epochs per task, type of non-stationarity, and choice of regularization strength; and (iii) achieves better generalization performance over the course of continual learning.
Datasets, Nonstationarities, and Architectures
Our experiments use the following datasets for continual supervised learning: MNIST [45], Fashion MNIST [46], EMNIST [47], SVHN2 [26], CIFAR10, and CIFAR100 [27]. In addition to the dataset, we consider three different types of non-stationarity: (i) random label assignments, (ii) pixel permutation, and (iii) label flipping. Random label assignments are commonly used to evaluate trainability [5] due to the large distribution shift between tasks in memorizing completely new and random labels. Pixel permutations, on the other hand, require only learning the permutation mask applied to the image; it induces loss of trainability more slowly but is useful for evaluating generalization [17]. Lastly, label flipping is a re-assignment of all the observations from one label to another label. Unlike the other two non-stationarities, a label flip distribution shift does not require learning a new representation. This is because only the output layer of a neural network needs to be permuted to learn the label re-assignment, but gradients can still unnecessarily change the representation leading to loss of trainability [12]. For the experiments on MNIST, Fashion MNIST and EMNIST, we use feed-forward neural networks. On SVHN2, CIFAR10, and CIFAR100 we use both a ResNet-18 [25] and Vision Transformer [24].
Loss of Trainability Mitigators
In our main results, we compare Wasserstein regularization and spectral regularization against L2 regularization towards zero, L2 regularization towards the initialization [17], recycling dormant neurons [23], shrink and perturb [15] and no regularization. In Appendix C.1, we provide additional results on other baselines: AdamW, which also regularizes towards zero and concatenated ReLU [4, 48]. Because we are interested in trainability in the task-agnostic setting, we do not compare with regularizers for catastrophic forgetting as they rely on privileged information regarding the task [49, 50]. Our experiments largely use Layer Norm [44] or Batch Norm in the case of ResNet-18 [51], which has been shown to be a mitigator for loss of trainability [37]. We also provide an ablation study showing that Layer Norm is synergistic with the regularizers we propose.
5.1 Comparative Evaluation
Figure 1: Comparison of task accuracy with random label assignment. Spectral regularization is particularly effective at mitigating loss of trainability. Without regularization, trainability is always reduced.
In Figure 1, we plot the results of training a ResNet-18 on different image classification datasets with random label assignment as the type of non-stationarity for each task. Across all problems and methods considered, spectral regularization was best at improving and sustaining its trainability, with accuracy increasing as new tasks were encountered. In contrast, despite improved trainability over the first few tasks, the unregularized baseline eventually experienced loss of trainability, leading to lower accuracy. Recycling dormant neurons, which reinitializes weights based on dormant neurons, experienced loss of trainability on all problems. One potential reason for this is because this approach only resets the weights of the convolutional layers and the final fully-connected layer [23], whereas our approach regularizes all parameters, including bias terms and batch normalization parameters. In Appendix C.1, we show that weight reinitialization can be effective on smaller problems without normalization layers. Shrink and perturb is similar to L2 regularization towards zero, but injects noise and is not as effective. While the perturbations were previously demonstrated to improve generalization when warm-starting a neural network on a new problem, we see that they are not effective when training on a longer sequence of problems. Wasserstein regularization is often on par with L2 regularization towards the initialization, and surpasses it on CIFAR10. These results demonstrate that, in contrast to other approaches, spectral regularization consistently maintains trainability across problems.
5.2 Sensitivity Analysis
In this section, we investigate the robustness of spectral regularization when varying (i) the number of training epochs per task, (ii) the type of non-stationarity applied, or, (iii) the regularization strength.
Varying the Number of Epochs Per Task
Given that the continual learning problem outlined in Section 2 depends on the number of iterations, it may be the case that a higher number of epochs per task can mitigate loss of trainability. In Figure 2, we found that the number of epochs only delays loss of trainability (see Appendix C.3 for results on Fashion MNIST). Even when the number of epochs per task is high enough to reach 100% accuracy on the first task, loss of trainability eventually occurred without regularization. In contrast, when using spectral regularization, loss of trainabililty was consistently mitigated. We additionally found that spectral regularization is particularly effective at maintaining trainability when the number of epochs per tasks is low. The evidence for this is most striking without Layer Norm, in the top row, where spectral regularization was the only method capable of maintaining its trainability. Even though Layer Norm controls the spectral norm through the activations [52], adding Layer Norm alone was not enough to mitigate loss of trainability. However, Layer Norm is synergistic with various forms of regularization, with accuracy being improved regardless of the number of epochs per task.
Figure 2: Sensitivity to number of epochs per task on MNIST with random label assignment. Without Layer Norm, spectral regularization is able to maintain trainability over other regularizers even when the number of epochs is low. Spectral regularization also synergizes with Layer Norm, improving initial performance on the first task, and improving over tasks.
Figure 3: Sensitivity to type of non-stationarity on EMNIST. Both Wasserstein and spectral regularization are able to consistently maintain high trainability across random label assignment, pixel permutations, and label flipping. When applying random label assignment, L2 (init) and L2 (zero) are unable to attain high trainability as compared to Wasserstein and spectral regularization.
Varying Type of Non-stationarity
Next, we explore whether our proposed regularizers are effective across different types of non-stationarity. To this end, we use EMNIST because the number of classes is large enough that label flipping induces loss of trainability [12]. In Figure 3, we found that both Wasserstein and spectral regularization best maintained trainability across different non-stationarities.
Figure 4: Sensitivity to regularization strength on MNIST with random label assignment. Spectral regularization is insensitive to regularization strength while remaining performant at few epochs.
Varying Regularization Strength
The effectiveness of a regularizer depends on the regularization strength. Too much regularization can slow down training, leading to suboptimal performance on any given task. However, too little regularization can lead to loss of trainability and suboptimal performance on future tasks. In a continual learning problem, hyperparameter tuning is particularly costly (or infeasible), especially if the distribution over tasks is not known in advance [53]. Thus, we should favour regularizers that are less sensitive to hyperparameters. Our results in Figure 4 show that spectral regularization dominates other regularizers for any choice of regularization strength, is much highly insensitive to the choice of hyperparameter compared to other regularizers.
Figure 5: Effects of regularization on distance from initialization, average condition, average weight rank, and gradient diversity when training ResNet-18 on SVHN2 with random label assignment. Despite being far from initialization, spectral regularization has a low average condition number, a high effective rank across its parameters, and a high effective gradient diversity.
5.3 Looking Inside the Network
We now explore structural properties of the neural network, and how these properties are effected by our proposed regularizers. In Figure 5, we further confirm that spectral regularization is the least restrictive approach, allowing deviations far from the initialization. Spectral regularization also controlled the spectral norm of the network, leading to a better average condition number (the ratio between the maximum and the minimum singular values), a higher effective rank of its parameters, and a higher gradient diversity. See Appendix C.5 for similar results on the other datasets considered.
Figure 6: Comparison of generalization performance with random label assignment. Spectral regularization improves generalization, using either a ResNet-18 (Top) or a Vision Transformer (bottom).
5.4 From Trainability to Generalization
Having demonstrated the main goal of this paper, namely the effectiveness of Wasserstein and spectral regularization on trainability, we now investigate generalization. We used a pixel permutation non-stationarity in Figure 6, and found that spectral regularization is better able to generalize compared to other regularizers, using either a ResNet-18 or Vision Transfomer. Interestingly, we also find that the Vision Transformer, particularly with spectral regularization, provides better generalization performance than using a ResNet. This demonstrates that regularizers which more directly address the issues underlying trainability also lead to better generalization.
6 Conclusion
In this paper, we investigated the connection between initialization, trainability and regularization. We identified that controlling the singular values so that they do not deviate significantly from their values at initialization promotes diverse gradients for neural networks that sustains trainability throughout continual learning. As the first approach to preserve trainability properties induced by initialization, building on recent work, we proposed to regularize the distribution of individual parameters during training towards the distribution at initialization. We then proposed to use spectral regularization as a way of controlling the spectral norm of each layer, thereby sustaining the trainability properties that are present at initialization. Our experiments show that spectral regularization is more performant and less sensitive to hyperparameters than other regularizers across datasets, nonstationarities, and architectures. While several methods mitigate loss of trainability to a varying degree, learning continually with spectral regularization is robust, achieving high initial and sustained performance. Finally, we show that spectral regularization is also capable of improving generalization with both Vision Transformer and ResNet-18 on continual versions of SVHN2, CIFAR10 and CIFAR100.
Appendix A Additional Details
A.1 Common Assumptions Regarding Tasks
We denote the parameters of a randomly initialized neural network trained only on task $\tau$ for $T$ iterations as $\theta_{\tau}^{(T)}=\theta^{(T-1)}_{\tau}-\alpha\nabla_{\theta}J_{\tau}(\theta)\big{|}_{\theta=\theta_{\tau}^{(t-1)}}$, with $\theta^{(0)}_{\tau}$ being sampled from a standard initialization distribution, such as Glorot [18] or He initialization [29]. The equal-difficulty assumption can be formalized by assuming that the objective value, $J_{\tau}(\theta_{\tau}^{(T)})$, is statistically independent of $\tau$. That is, while the data depends on the task, the objective value after training does not depend on the task. Note that the equal-difficulty assumption is key: otherwise it would be difficult to disambiguate whether a performance decrease is due to a more difficult task or due to loss of trainability in the learning algorithm.
A.2 Illustrative Example: Large Spectral Norm Can Impede Trainability
We show that, given a family of solutions on the first task, preference should be given to solutions that have a spectral norm that is closer to one. We consider 2 dimensional binary and orthogonal inputs, $x_{1}=[0,1]^{\top}$ and $x_{2}=[1,0]^{\top}$, with corresponding binary targets, $y_{1}=y_{2}=1$. We consider the squared loss, and use a mutli-layer perceptron with one hidden layer and two units, $f_{w_{2},w_{1}}(x)=w_{2}\texttt{ReLU}(w_{1}x)$. A common observation regarding solutions found for deep neural networks is that they are low rank [54, 55, 56, 57]. In addition, it has been observed that large outliers in the spectrum of a deep network emerge during training, meaning that the largest singular values of the weight matrices tend to be much larger than their smallest singular values [58, 22]. Consider one such solution where $w_{1}=\begin{bmatrix}-1&c\\ \nicefrac{{1}}{{c}}&-1\end{bmatrix}$, and $w_{2}=[\nicefrac{{1}}{{c}},c]\in\mathbb{R}^{1\times 2}$. The input-output function is invariant under different values of $c$, achieving zero error for any particular choice. However, the parameter matrix for the hidden activations, $w_{1}$, is low rank because it only has a single non-zero singular value that is not invariant. That is, the largest singular value of $w_{1}$ depends on the choice of $c$, $\|w_{1}\|_{2}:=\sigma_{max}(w_{1})=c+\nicefrac{{1}}{{c}}$. In addition, the gradients and parameter updates depend on $c$. Thus, if the targets were to change, $y_{1}^{\prime}=0$, the imbalance of singular values due to very large or small values of $c$ can impede trainability:
$w_{1}^{\prime}(x_{1})=\begin{bmatrix}-1&c\\ \nicefrac{{1}}{{c}}&-1\end{bmatrix}-\alpha\left(\underbrace{\begin{bmatrix}0&\nicefrac{{1}}{{c}}\\ 0&0\end{bmatrix}}_{\nabla_{w_{1}}\ell_{1}}\right),\hskip 28.45274ptw_{2}^{\prime}(x_{1})=[\nicefrac{{1}}{{c}},c]-\alpha\left(\underbrace{[c,0]}_{\nabla_{w_{2}}\ell_{1}}\right),$ (1)
where $\ell_{1}=\ell(f_{w_{2},w_{1}}(x_{1}),y^{\prime}_{1})$, and $\alpha$ is a step-size. One of the consequences of the imbalanced singular values is that the per-example gradients are poorly conditioned [59]. That is, when we sample $x_{1}$, there will be a large gradient for $w_{2}$, $\|\nabla_{w_{2}}\ell(f_{w_{2},w_{1}}(x_{1}),y_{1}^{\prime})\|^{2}=c$ but a small gradient for $w_{1}$, $\|\nabla_{w_{1}}\ell(f_{w_{2},w_{1}}(x_{1}),y_{1}^{\prime})\|^{2}=\nicefrac{{1}}{{c}}$. Thus, a sufficiently small step-size that stabilizes the update to $w_{2}$ leads to slow learning on $w_{1}$. A similar result holds if the target $y_{2}$ were to change In Figure 7, we demonstrate that learning is more stable with a smaller spectral norm in this example. For higher dimensional weight matrices, we may expect that the imbalance of the singular values leads to low gradient diversity where particular per-example gradients give rise to outliers with a large gradient magnitudes. Thus, all else being equal, a solution with a smaller spectral norm ($c\approx 1$) is preferable for the purposes of continual learning.
Figure 7: Illustrative Example. A lower spectral norm is more trainable across different step sizes in the simple illustrative example.
A.3 Differences Between Continual and Non-continual Learning
One critical difference between continual and non-continual learning is that, in continual learning, convergence towards a fixed point is not the objective. Due to continual changes in the data distribution, there is no fixed point that is optimal for every distribution in general. Thus, convergence to a fixed point is necessarily suboptimal. If convergence to a fixed point does occur, then the norm of the gradients will be zero. The gradients converging to zero is a condition for loss of plasticity, and can be satisfied on every distribution depending on the choice of activation function. Focusing on the case when the activation function is $\texttt{ReLU}(x)=\max(x,0)$, the gradients can become zero if 1) the weights converge to zero or, 2) the activations converge to zero. As a running example, we will consider the two-layer neural network without biases denoted by $f_{\theta}(x)=\theta_{2}\phi(\theta_{1}x)$, where $\theta_{1}\in\mathbb{R}^{d\times d_{in}}$ and $\theta_{2}\in\mathbb{R}^{d_{out}\times d}$ are weight matrices, and $\phi$ is an elementwise activation function. We will consider the squared loss, $J(\theta)=\frac{1}{2}\mathbb{E}_{(x,y)\sim p}\left[(y-f_{\theta}(x))^{2}\right]$, but a similar result holds for classification losses, such as cross-entropy. The gradient for each of the parameters is given by, $\nabla_{\theta_{2}}J(\theta)=(y-f)\phi(\theta_{1}x)^{\intercal}$ and $\nabla_{\theta_{1}}J(\theta)=\theta_{2}^{\intercal}(f-y)\phi^{\prime}(\theta_{1}x)x^{\intercal}$. The neural network will reach a fixed point in continual learning, meaning that the gradients are zero for every datapoint, if either 1) both $\theta_{1}$ and $\theta_{2}$ are zero, or 2) the activations are zero for every input $x$, $\phi(\theta_{1}x)=0$,
A.4 Spectral Norm and Lipschitz Constant
One other potential reason for why a growing spectral norm is problematic is from the theory convex optimisation. For the ReLU activation function, the Lipschitz constant of the layer map, $h_{l+1}=\phi(\theta_{l}h_{l})$, is equal to the spectral norm of the weight matrix. We can thus bound the Lipschitz constant of the entire neural network above by the product of the layer-wise Lipschitz constants [60]. Putting these results together, we conclude that during continual learning, the Lipschitz constant of each layer is increasing and thus the Lipschitz constant of the entire network is also increasing. This is problematic from the standpoint of optimisation because gradient descent only converges on convex optimisation problems when the step-size is smaller than $\frac{2}{L}$, where $L$ is the Lipschitz constant. Thus, if the Lipschitz constant grows then, eventually, the chosen step-size will be too large and gradient descent will not converge locally on the task.
A.5 More Details on Regularization and Trainability
One other important advantage of regularization is that it has well understood effects for ensuring trainability in approaches outside of deep learning, such as degenerate linear regression and inverse problems [61].
Each regularizer biases the parameter dynamics in their own way, such as keeping the parameters close to zero or close to initialization. But, the regularizer term is independent of the base objective, and potentially the data and/or architecture. Even if the base objective provides a zero gradient, explicit regularization can provide a gradient to all weights, and reset weights if the hidden unit associated with the weight becomes saturated. For example, if the nonlinearity is ReLU, then a hidden unit ($i$) is inactive if for all inputs in the dataset ($k$), $h^{l}_{i}=\phi\left(\sum_{j=1}^{d_{l-1}}\theta^{l}_{ij}h^{l-1}_{j,k}\right)=0$. Then the hidden representation has collapsed for that unit, and none of the weights contributing to that unit will be updated, keeping the unit inactive. But, with regularization, the weights will be updated and the unit may become active again for some inputs.
A.6 More Details on Wasserstein Regularizer
At initialization, the parameters at layer $l$ are sampled i.i.d. $\mathbf{\theta}_{i,j}\sim p^{(l,0)}(\theta)$ according to some pre-determined distribution, such as the Glorot initialization [18]. The distribution of parameters at iteration $t$ during training and for any particular layer, denoted by $p^{(l,t)}$, is no longer known and does not have a closed form. However, it is still possible to regularize the empirical distribution towards the initialization distribution by using the empirical Wasserstein metric [39]. We denote the flattened parameter matrix for layer $l$ at time $t$ by $\mathbf{\bar{\theta}}^{(l,t)}$. The squared Wasserstein-2 distance between the distribution of parameters at initialization and the current parameter distribution is defined as
$\mathcal{R}_{\tau(t)}(\theta,\theta^{(0)})=\sum_{i=1}^{d}\left(\bar{\theta}_{(i)}^{(l,t)}-\bar{\theta}_{(i)}^{(l,0)}\right)^{2}.$
Unlike the regenerative regularizer, the Wasserstein regularizer takes the difference of the order statistics. Thus, the regenerative regularizer is always larger because the Wasserstein regularizer takes the difference in the sorted values, $\sum_{i=1}^{d}\left(\mathbf{\bar{\theta}}_{(i)}^{(l,t)}-\mathbf{\bar{\theta}}_{(i)}^{(l,0)}\right)^{2}\leq\sum_{i=1}^{d}\left(\mathbf{\bar{\theta}}_{i}^{(l,t)}-\mathbf{\bar{\theta}}_{i}^{(l,0)}\right)^{2}$. The Wasserstein regularizer, like variants of L2, regularize the individual parameters and may not preserve the properties of the entire parameter matrix at each layer.
A.7 Details Regarding Spectral Regularization
Spectral Regularization vs L2 Regularization
L2 regularization is not as effective as spectral regularization because it constrains the magnitude of the parameters, and can even cause rank collapse [17]. Spectral regularization constrains only the spectral norm of the weight matrix close to 1, $\sigma^{(1)}(\theta_{l})\approx 1$. This has the effect of controlling the magnitude of the maximal column of the weight matrix. That is, if $\theta\in\mathbb{R}^{d_{l}\times d_{l-1}}$, then $\nicefrac{{1}}{{\sqrt{d_{l}}}}\|\theta_{l}\|_{1}<\|\theta_{l}\|_{2}$ where $\|\theta_{l}\|_{1}=\max_{1\leq j\leq d_{l}}\sum_{i=1}^{d_{l}}|[\theta_{l}]_{ij}|$. This implies that $\sum_{i=1}^{d_{l}}|[\theta_{l}]_{ij}|<\sqrt{d_{l}}$ for every column $j$. This differs in two ways from L2 regularization: (i) the regularizer does not regularize all the columns of the weight matrix, only the parameters in the column with a maximal sum, and (ii) the regularizer does not regularize the individual weights of the column, but their sum. This means that the parameters can potentially be larger, and move further from the initialization, potentially enabling more effective use of the neural network’s capacity. Lastly, we note that L2 regularization can also be viewed in terms of a matrix norm, specifically the Frobenius norm.
Spectral Regularization vs Spectral Normalization:
The spectral norm can also be controlled via spectral normalization [43]. For continual learning, spectral regularization is preferable over spectral normalization for two reasons: (i) spectral normalization is data-dependent (see Section 2.1 and Equation 12 in [43], and Appendix A.8 for more details), which can be problematic in continual learning due to the changing data distribution, and, (ii) other forms of normalization are often already used to train deep neural networks, such as LayerNorm [44]. LayerNorm, in particular, regularizes the magnitude of the layerwise map by dividing by the norm of the features in the layer. However, because the spectral norm of the underlying parameters still grows at a rate of $\sqrt{t}$, the normalization layer itself will suffer from loss of gradient diversity. For the purposes of ensuring continual trainability, regularization explicitly controls the spectral norm of all parameters and is thus preferable over normalization in isolation. That being said, normalization can still improve performance and optimal performance may be achieved by a combination of regularization and normalization, which we show in Section 5.2.
A.8 Categorizing Regularizers for Continual Learning
We can categorize regularizers broadly into data-dependent and data-indepedent regularizers. Data-dependent regularizers use the data in some way, which can be probelmatic for continual learning when the data distribution changes. Regularizing on one distribution of data does not necessarily maintain useful properties for learning on the next distribution of data. Examples of data-dependent regularizers include feature rank regularization [62], auxiliary tasks and other feature-space regularizers [14]. Other non-standard approaches that have a data-dependent regularization effect include gradient regularizers such as gradient-clipping and weight reinitialization based on dormant neurons [23]. These data-dependent regularizers regularizers control some property on data from the current task but may not control the property on data from a new task, which is needed to maintain plasticity.
Data-Independent Regularization, on the other hand, does not depend on any data. This category regularizes the parameters directly, which is particularly useful in continual learning when the data distribution is changing. Examples of data-independent regularization include L2 regularization towards zero, weight decay (for adaptive gradient methods), and regenerative regularization [17]. Our proposed approach of spectral regularization is data-independent and, moreover, explicitly targets a trainability condition for continual learning. Thus, we expect that it is particularly effective on maintaining trainability in continual learning compared to other data-independent regularizers.
A.9 Spectral Regularization of Other Layers
Normalization Layers
Other layers using per-unit scaling, such as normalization layers [51, 44], are also spectrally regularized. For example, denote $\gamma\in\mathbb{R}^{d}$ as the trainable scaling parameters for LayerNorm [44], then the element-wise product of the weights can be written as a diagonal weight matrix, $f(z)=\gamma\circ z=\text{Diag}(\gamma)z$. The maximum singular value of a diagonal matrix is the maximum entry, $\sigma_{max}(\text{Diag}(\gamma))=\max_{i}|\gamma_{i}|$. However, optimisation with a maximum is problematic because the max is not differentiable. Furthermore, differentiable surrogates, such as log-sum-exp ($\left(\log\left(\sum_{i}\exp(|\gamma_{i}|)\right)-1\right)^{2}$), only give a loose upper bound on the maximum which is problematic because we want to regularize the maximum towards one. Thus, we regularize each weight towards 1.
Convolutional Layers
Similar to other work, we reshape the convolutional weight tensor (with kernel size ($k\times k$, $d_{in}$ filters and $d_{out}$ filters) into a matrix of size $d_{out}\times(k\cdot k\cdot d_{in})$ [42]. The spectral norm of this reshaped matrix provides an efficient upper bound on the spectral norm of the Toeplitz matrix defining the convolution [63, Corollary 1].
Appendix B Experiment Details
All of our experiments used Adam [64] where the default step size of $0.001$ was selected after an initial sweep over $[0.005,0.001,0.0005]$. For all of our results, we use 10 random seeds and provided a shaded region corresponding to the standard error of the mean. For experiments on SVHN2, CIFAR10 and CIFAR100, we used 5 seeds to sweep over the regularization strengths of $[0.01,0.001,0.0001]$, and found that $0.0001$ worked well on CIFAR10 and CIFAR100 for all regularizers, whereas $0.001$ worked best on SVHN2 for all regularizers.
Datasets and non-stationarities:
  • MNIST: Only the first 12800 datapoints were used for training, with a batch size of 512 and [40, 80, 120] epochs, and a total of 50 tasks.
  • Fashion MNIST: Only the first 12800 images, with a batch size of 512 and [40, 80, 120] epochs per task, and a total of 50 tasks.
  • EMNIST: We use the balanced version of the dataset, using the first 100000 datapoints. For random label non-stationarity, we used a batch size of 500 and 100 epochs per task, with a total of 50 tasks. For both label flipping and pixel permutation non-stationarities, we used a batch size of 500 and 20 epochs per task, with a total of 200 tasks.
  • SVHN2: The first 50000 images were used for training, 5000 images from the test set were used for validation and the rest were used for testing. The batch size used was 500, we found 250 to be unreliable for learning due to high variance. For random label non-stationarity, 20 epochs per task and 25 tasks was enough to lose trainability for an unregularized network. For pixel permutation non-stationarity, 10 epochs and 100 tasks was enough to lose trainability for an unregularized network.
  • CIFAR10: All of the 50000 images were used for training, 1000 images from the test set were used for validation and the rest were used for testing. We used a batch size of 250 and found this to be effective. For random label non-stationarity, 20 epochs per task and 25 tasks was enough to lose trainability for an unregularized network. For pixel permutation non-stationarity, 10 epochs and 100 tasks was enough to lose trainability for an unregularized network.
  • CIFAR100: All of the 50000 images were used for training, 1000 images from the test set were used for validation and the rest were used for testing. We used a batch size of 250 and found this to be effective. For random label non-stationarity, 20 epochs per task and 25 tasks was enough to lose trainability for an unregularized network. For pixel permutation non-stationarity, 10 epochs and 100 tasks was enough to lose trainability for an unregularized network.
Neural Network Architectures:
  • For MNIST, EMNIST, and Fashion MNIST: 4-layer MLP with 256 neurons per layer and relu activations. For experiments that use it, LayerNorm was applied after the linear weight matrix and before the non-linearity.
  • For SVHN2, CIFAR10, and CIFAR100: An off-the-shelf ResNet-18 with batch norm, as well as an off-the-shelf Vision Transformer (tiny).
Appendix C Additional Experiments
C.1 Other Baselines
Figure 8: AdamW in Random Label MNIST: AdamW performs similarly to L2 regularization towards zero, recovering trainability when LayerNorm is used.
Figure 9: Concatenated ReLU, Random Label MNIST. CReLU is unable to maintain trainability even with layer normalization.
Figure 10: Recycling Dormant Neuron, Random Label MNIST. Recycling dormant neurons are effective without layer normalization, but layer normalization removes dormant neurons and still leads to loss of trainability.
C.2 Ablating Hyperparameter $k$ in Spectral Regularization
Figure 11: Evaluating the choice of $k$ for spectral regularization for $k=1,2,4,8$ (Left) MNIST, (Right) Fashion MNIST.
C.3 Results on Fashion MNIST
Figure 12: Loss of Plasticity in Fashion MNIST Regularization is necessary, LayerNorm without regularization loses plasticity: the performance difference increases as a function of tasks, even with an increasing number of iterations.
C.4 Effect of Regularization on Capacity in a Single Task
The regularization strength for preventing loss of trainability must be sufficiently high, and this regularization strength may limit the capacity of the neural network. We now show the extended training of each regularizer using the same regularization strength used to prevent loss of trainability in Figure 13. Spectral regularization was not only best in preventing loss of trainability, it also achieved the highest accuracy when training to convergence on a single task. This means that spectral regularization constrains the capacity of the neural network the least, while still preventing loss of trainability.
Figure 13: Single task performance with ResNet-18. Spectral regularization is least restrictive of the neural network capacity, evidenced by its ability to better fit the randomly assigned labels compared to other regularizers. The unregularized baseline is able to use its capacity fully, but at the cost of reduced trainability in later tasks.
C.5 Investigating Effects of Regularization on Neural Network Properties
Figure 14: Singular Values and Weight Norms: Top: SVHN2, Middle: CIFAR10, Bottom: CIFAR100.
Figure 15: Rank and Singular Value Spread: Top: SVHN2, Middle: CIFAR10, Bottom: CIFAR100.
C.6 Investigating Generalization on Continual Imagenet
We investigate the generalization performance of spectral regularization on Continual ImageNet. In Continual ImageNet, each task is to distinguish between two ImageNet classes. We use the network architecture and training protocol used in [17]. We find that all regularization approaches tested achieve high generalization performance across tasks. The methods we compare are spectral regularization (SpectralRegAgent), L2 towards zero (L2Agent), L2 towards initialization (L2InitAgent), recycling dormant neurons (ReDOAgent), and no regularization (BaseAgent).
Figure 16: Generalization performance on Continual ImageNet.

### v2

@@
 Abstract
-footnotetext: Department of Computing Science, University of Alberta
-Department of Computer Science, Stanford University
-Google DeepMind
-Canada CIFAR AI Chair
-Loss of plasticity is a phenomenon where neural networks become more difficult to train during the course of learning.
-Continual learning algorithms seek to mitigate this effect by sustaining good predictive performance while maintaining network trainability.
-We develop new techniques for improving continual learning by first reconsidering how initialization can ensure trainability during early phases of learning.
-From this perspective, we derive new regularization strategies for continual learning that ensure beneficial initialization properties are better maintained throughout training.
-In particular, we investigate two new regularization techniques for continual learning: (i) Wasserstein regularization toward the initial weight distribution, which is less restrictive than regularizing toward initial weights; and (ii) regularizing weight matrix singular values, which directly ensures gradient diversity is maintained throughout training.
-We present an experimental analysis that shows these alternative regularizers can improve continual learning performance across a range of supervised learning tasks and model architectures.
-The alternative regularizers prove to be less sensitive to hyperparameters while demonstrating better training in individual tasks, sustaining trainability as new tasks arrive, and achieving better generalization performance.
+footnotetext: Correspondence to: Alex Lewandowski <lewandowski@ualberta.ca>.
+Loss of plasticity is a phenomenon where neural networks can become more difficult to train over the course of learning.
+Continual learning algorithms seek to mitigate this effect by sustaining good performance while maintaining network trainability.
+We develop a new technique for improving continual learning inspired by the observation that the singular values of the neural network parameters at initialization are an important factor for trainability during early phases of learning.
+From this perspective, we derive a new spectral regularizer for continual learning that better sustains these beneficial initialization properties throughout training.
+In particular, the regularizer keeps the maximum singular value of each layer close to one.
+Spectral regularization directly ensures that gradient diversity is maintained throughout training, which promotes continual trainability, while minimally interfering with performance in a single task.
+We present an experimental analysis that shows how the proposed spectral regularizer can sustain trainability and performance across a range of model architectures in continual supervised and reinforcement learning settings.
+Spectral regularization is less sensitive to hyperparameters while demonstrating better training in individual tasks, sustaining trainability as new tasks arrive, and achieving better generalization performance.
 1 Introduction
-A longstanding goal of machine learning research is to develop algorithms that can learn continually and cope with unforeseen changes in the data distribution [1, 2].
-Current learning algorithms, however, struggle to learn from dynamically changing targets and are unable to adapt gracefully to unforeseen changes in the distribution during the learning process [3, 4, 5, 6].
-Such limitations can be seen as a byproduct of assuming, one way or another, the problem is stationary.
-Recently, there has been growing recognition of the fact that there are limitations to what can be learned from a fixed and unchanging dataset [7], that there are implicit non-stationarities in many problems of interest [8], and that some real-world problems benefit from learning continually [9, 10].
-The concept of plasticity has been receiving growing attention in the continual learning literature, where the loss of plasticity—a reduction in the trainability of the neural network [11, 5, 12]—has been noted as a critical shortcoming in current learning algorithms.
-Learning algorithms, and more specifically neural networks, that are performant in the non-continual learning setting often struggle when applied to continual learning problems.
-Settings where a neural network must continue to learn after changes occur in the observations exhibit a striking loss of plasticity such that learning slows down or even halts after successive changes in the learning distribution [13, 6, 14].
-Several aspects of a learning algorithm have been found to contribute to, or mitigate, loss of plasticity.
-Examples include the type of optimizer [6], the step-size [15, 16], the number of optimiser iterations [5], and the use of specific regularizers [11, 14, 17].
+A longstanding goal of machine learning research is to develop algorithms that can learn continually and cope with unforeseen changes in the data distribution (Ring, 1994, Thrun, 1998).
+Current learning algorithms, however, struggle to learn from dynamically changing targets and are unable to adapt gracefully to unforeseen changes in the distribution during the learning process (Zilly et al., 2021, Abbas et al., 2023, Lyle et al., 2023, Dohare et al., 2024).
+Such limitations can be seen as a byproduct of assuming, one way or another, that the problem is stationary.
+Recently, there has been growing recognition of the fact that there are limitations to what can be learned from a fixed and unchanging dataset (Hoffmann et al., 2022), that there are implicit non-stationarities in many problems of interest (Igl et al., 2021), and that some real-world problems benefit from learning continually (Han et al., 2022, Janjua et al., 2023).
+The concept of plasticity has been receiving growing attention in the continual learning literature, where the loss of plasticity—either a reduction in a neural network’s ability to train (Dohare et al., 2021, Lyle et al., 2022, Elsayed and Mahmood, 2024), or ability to generalize (Ash and Adams, 2020, Zilly et al., 2021)—has been noted as a critical shortcoming in current learning algorithms.
+Learning algorithms, and more specifically neural networks, that are performant in the non-continual learning setting, often struggle when applied to continual learning problems.
+Settings where a neural network must continue to learn after changes occur in the data distribution exhibit a striking loss of plasticity such that learning slows down (Lyle et al., 2023) or even halts after successive changes (Abbas et al., 2023, Nikishin et al., 2022, Dohare et al., 2024).
+Several aspects of a learning algorithm have been found to contribute to or mitigate loss of plasticity.
+Examples include the type of optimizer (Dohare et al., 2024, Lyle et al., 2023), the step-size (Ash and Adams, 2020, Berariu et al., 2021), the number of optimiser iterations (Lyle et al., 2023), and the use of specific regularizers (Dohare et al., 2021, Lyle et al., 2022, Kumar et al., 2023, Lewandowski et al., 2023).
 Such factors hint that there might be simpler underlying optimisation principles that govern the loss of plasticity.
-For example, the success of several methods that regularize neural networks towards properties of the initialization suggests that some of those properties mitigate loss of plasticity [11, 17, 14].
-Various properties of neural network weights at initialization have been shown to be associated with trainability.
-Previous analyses have demonstrated that the trainability of deep neural networks can be improved by ensuring that the variance of hidden activations and gradients remains uniform across the different layers of the neural network [18], which has the effect of keeping the average singular value of the layerwise Jacobians close to one.
-A stronger condition is dynamical isometry, where all singular values are close to one, or exactly one in the case of an orthogonal initialization [19, 20, 21].
-Over the course of learning, the singular values deviate from their initialization and tend to grow over time [22], which can reduce trainability and impede continual learning.
+For example, the success of several methods that regularize neural networks towards properties of the initialization suggests that some of those properties mitigate loss of plasticity (Dohare et al., 2021, Kumar et al., 2023, Lyle et al., 2022).
+The properties of neural network weights at initialization are associated with trainability in the early phases of learning.
+Previous analyses have demonstrated that the trainability of deep neural networks can be improved by ensuring that the initialization variance of hidden activations and gradients remains uniform across the different layers of the neural network (Glorot and Bengio, 2010), which has the effect of keeping the average singular value of the layerwise Jacobians close to one.
+A stronger condition is dynamical isometry, where all singular values are close to one, or exactly one in the case of an orthogonal initialization (Saxe et al., 2014, Pennington et al., 2017, Xiao et al., 2018).
+Over the course of learning, the singular values deviate from their initialization and tend to grow over time (Martin and Mahoney, 2021), which can reduce trainability and impede continual learning.
 Given the effectiveness of initialization in ensuring the trainability of a neural network, continual learning algorithms may benefit from sustaining the relevant properties of the weights.
-This is the motivation of several methods that use the initialization either through regularization [17] or weight reinitialization [11, 23, 6].
-Such methods leverage initialization explicitly, either by regularizing parameters towards their initial values or by resetting parameters by resampling them from the initialization distribution.
-In this paper, we seek to more directly address loss of trainability by sustaining the key properties present at initialization, rather than replicate the initialization explicitly.
+This is the motivation of several methods that use the initialization either through regularization (Kumar et al., 2023) or weight reinitialization (Dohare et al., 2021, Sokar et al., 2023, Dohare et al., 2024).
+Such methods leverage the initialization explicitly, either by regularizing parameters towards their initial values or by resetting parameters by resampling them from the initialization distribution.
+In this paper, we seek to more directly address the loss of trainability by sustaining the key properties present at initialization, thereby striking a better balance between trainability and performance compared to approaches that replicate the initialization explicitly.
 We investigate the key properties of the initialization that ensure trainability, how these properties are lost over the course of learning, and the effects this has on continual learning.
-Based on this analysis, we then introduce two regularizers that try to maintain the identified properties: we propose to use the Wasserstein distance to regularize the parameters towards the initialization distribution, which is less restrictive than regularizing towards initial weights, and we propose spectral regularization to control the maximum singular value of the weight matrix to directly address the deviation of the singular values from initialization over the course of learning.
+We identify that deviations from the initial singular value distribution can result in low gradient diversity, thereby impeding continual learning.
+Based on this analysis, we then introduce a spectral regularizer to control the deviation of the singular values by keeping the maximum singular value of each layer close to one to directly address the deviation of the singular values from initialization over the course of learning.
 Our experiments show that spectral regularization is more performant and less sensitive to hyperparameters than other regularizers across datasets, nonstationarities, and architectures.
 While well-tuned regularizers are often able to mitigate loss of plasticity to a varying degree, learning continually with spectral regularization is robust, achieving high initial and sustained performance.
-Finally, we show that spectral regularization is also capable of improving generalization with both Vision Transformer [24] and ResNet-18 [25] on continual versions of SVHN2 [26], CIFAR10, and CIFAR100 [27].
+In particular, we show that spectral regularization is also capable of improving generalization with both Vision Transformer (Dosovitskiy et al., 2021) and ResNet-18 (He et al., 2016) on continual versions of tiny-ImageNet (Le and Yang, 2015), CIFAR10, CIFAR100 (Krizhevsky, 2009), and SVHN2 (Netzer et al., 2011).
+Note that these datasets and architectures encompass all continual supervised learning experimental settings considered in the loss of plasticity literature.
+We also show that spectral regularization can improve the performance of soft-actor critic (Haarnoja et al., 2018) in reinforcement learning settings where loss of plasticity occurs due to primacy bias (Nikishin et al., 2022).
 2 Problem Setting
-We investigate the trainability of learning algorithms, and specifically neural networks, in task-agnostic continual supervised learning.
-We assume the learning algorithm uses a variant of stochastic gradient descent to update a neural network with $L$ layers, defined recursively as $f_{\theta}(x):=h_{L}(x)$, where $h_{0}(x)=x$ is the input vector, $h_{l+1}(x)=\phi(\mathbf{W}_{l}h_{l}(x)+\mathbf{b}_{l})$ with an element-wise activation function $\phi$ and parameters $\theta=\{\theta_{L},\dots,\theta_{1}\}$ where $\theta_{l}=\{\mathbf{W}_{l},\mathbf{b}_{l}\}$ is the weight and bias parameter for layer $l$ [^1] In this notation we suppressed the presence of biases, which can effectively be included in the parameters.
-We assume that $f_{\theta}$ is trained over a sequence of tasks: the $\tau$th task is specified by a distribution $p_{\tau}$ over the observation-target pairs (denoted by $(x,y)$).
-For simplicity we assume that the tasks (data distribution) change periodically after every $T$ iterations, but we consider the task-agnostic setting, where the learning algorithm does not have access to information about the tasks except through the data that it samples.
-For each iteration in the task, $t\in\left((\tau-1)T,\tau T\right]$, the learning algorithm must optimise the neural network’s parameters to minimize the objective
-$J_{\tau}(\theta)=\mathbb{E}_{(x,y)\sim p_{\tau}}\big{[}\ell(f_{\theta}(x),y)\big{]}$
+We investigate the trainability of neural network learning algorithms in the task-agnostic setting.
+We denote a neural network with $L$ layers, defined recursively as $f_{\theta}(x):=h_{L}(x)$, where $h_{0}(x)=x$ is the input vector, $h_{l+1}(x)=\phi(\mathbf{W}_{l}h_{l}(x)+\mathbf{b}_{l})$ with an element-wise activation function, $\phi$, and parameters $\theta=\{\theta_{L},\dots,\theta_{1}\}$ where $\theta_{l}=\{\mathbf{W}_{l},\mathbf{b}_{l}\}$ is the weight and bias parameter for layer $l$ [^1] In this notation we suppressed the presence of biases, which can effectively be included in the parameters.
+We assume that $f_{\theta}$ is trained over a sequence of tasks: the $\tau$th task is specified by a distribution $p_{\tau}$ over the observation-target pairs, denoted by $(x,y)$.
+For simplicity we assume that the task (data distribution) changes periodically after every $T$ iterations.
+In addition, we consider the task-agnostic setting, where the learning algorithm does not have access to information about the task except through the data that it samples.
+For each iteration in the task, $t\in\left((\tau-1)T,\tau T\right]$, the learning algorithm optimises the neural network’s parameters to minimize the objective
+$J_{\tau}(\theta)=\mathbb{E}_{(x,y)\sim p_{\tau}}\big{[}\ell(f_{\theta}(x),y)\big{]},$
 for some loss function $\ell$.
-In this paper we consider gradient-based methods to learn the weight parameters; the basic version of gradient descent would update the parameters at iteration $t$ as $\theta^{(t+1)}=\theta^{(t)}-\alpha\nabla_{\theta}J_{\tau(t)}(\theta)\big{|}_{\theta=\theta^{(t)}}$, where $\tau(t)=\lfloor t/T\rfloor$ denotes the current task number.
-In practice this is done by considering the empirical error instead of $J_{\tau(t)}$ computed from samples $(x_{t},y_{t})\sim p_{\tau}$ for $t\in\left((\tau-1)T,\tau T\right]$ generated from the task distribution, and usually some more involved optimization algorithm.
-The evaluation of the continual learning algorithm is performed at the end of a task [5].
+In this paper, we consider gradient-based methods to learn the weight parameters; the basic version of gradient descent would update the parameters at iteration $t$ as $\theta^{(t+1)}=\theta^{(t)}-\alpha\nabla_{\theta}J_{\tau(t)}(\theta)\big{|}_{\theta=\theta^{(t)}}$, where $\tau(t)=\lfloor t/T\rfloor$ denotes the current task number.
+In practice this is done by considering the empirical error instead of $J_{\tau(t)}$ using samples from the task distribution, $(x_{t},y_{t})\sim p_{\tau}$, and usually some more involved optimization algorithm.
+The evaluation of the continual learning algorithm is performed at the end of a task (Lyle et al., 2023).
 Specifically, the learning algorithm is evaluated at times $t=\tau T$, after being given $T$ iterations to learn on task $\tau$.
-It is commonly assumed that each task is sampled independently and identically, and that each task is equally difficult, meaning that a neural network which is randomly initialized on each task is able to reach a similar objective value for any particular task [6, 12, 5].
-Loss of plasticity in the continual learning literature can refer to either loss of trainability [11, 5] or to loss of generalization [15].
-Because trainability is a requirement for learning and generalization, we focus on loss of trainability, using it interchageably with loss of plasticity.
+A useful assumption in some problems is that each task is sampled independently and identically, and that each task is equally difficult, meaning that a neural network which is randomly initialized on each task is able to reach a similar objective value for any particular task (Lyle et al., 2023, Elsayed and Mahmood, 2024, Dohare et al., 2024).
+While this assumption is useful, it is not always applicable.
+In class-incremental learning (Van de Ven et al., 2022), the addition of more classes increases the difficulty of subsequent tasks.
+Another example is the transient tasks induced over the course of reinforcement learning (Igl et al., 2021), which can have varying task difficulty and lead to loss of plasticity due to primacy bias (Nikishin et al., 2022).
+We also consider such settings here.
+Loss of plasticity in the continual learning literature can refer to either loss of trainability (Dohare et al., 2021, Lyle et al., 2023) or to loss of generalization (Ash and Adams, 2020).
+Because trainability is a requirement for learning and generalization, we focus primarily on loss of trainability.
 Specifically, we use loss of trainability to refer to the phenomenon that the objective value, $J_{\tau}(\theta^{(\tau T)})$, increases as a function of the task $\tau$.
 Equivalently, the performance measures, such as accuracy, decrease with new tasks.
-Performance can also decrease, then plateau, and in this case the network has lost trainability.
+Under the assumption that the tasks are sampled independently and identically, this would suggest that the neural network’s trainability diminishes on new tasks.
 3 Spectral Properties and Continual Trainability
-A growing body of work has identified that neural networks trained on continually changing distributions exhibit loss of trainability, slowing down (or even preventing) learning on new tasks [11, 13, 6, 5].
-At a high level, the reason behind this phenomenon is simple: the solution of one task becomes the initialization for learning on the next task, and if this initialization is not sufficiently good to enable learning in the new task, we face the aforementioned problem.
+At a high level, the reason behind loss of plasticity is simple: the solution of one task becomes the initialization for learning on the next task, and if this initialization is not sufficiently good to enable learning in the new task, we face the aforementioned problem.
 To avoid this issue, we need to keep the network parameters within a region that can serve as good initialization at the start of a task.
-This approach comes with two challenges: (i) determining a suitable region for parameter initialization; (ii) ensuring the initialization region does not prevent effective learning in the tasks encountered.
-It is easy to satisfy (i) and (ii) in isolation: using a standard initialization algorithm, such as those of [28, 29], and then keeping the parameters fixed at this value satisfies (i) but clearly does not satisfy (ii).
+This approach comes with two challenges: (i) determining a suitable region for parameter initialization on any given task; (ii) ensuring that training within the initialization region does not prevent effective learning on the current task.
+It is easy to satisfy (i) and (ii) in isolation: using a standard initialization algorithm, such as those by Hinton and Salakhutdinov (2006), He et al.
+(2015), and then keeping the parameters fixed at this value satisfies (i) but clearly does not satisfy (ii).
 Conversely, not restricting the parameter space while using a standard learning algorithm addresses (ii) but not (i).
 We seek a balance of the two requirements.
-Note that (i) and (ii) are not binary properties, hence there can be a trade-off (e.g., allowing less favorable parameters for next task initialization allow better optimization in the current task).
-It is worth noting that such an approach has been successfully applied in the context of online convex optimization in changing environments, where several algorithms can be recognized as running a standard learning algorithm (such as mirror descent) on a carefully selected subset of the parameter space (see, e.g., [30, 31, 32]).
-Our approach is the following: to address problem (i), we first identify key properties that initialization algorithms [28, 29] impose on the starting parameters of a neural network.
-Then, we propose two regularizers targeting these properties strikes a good balance between (i) and (ii).
+Note that (i) and (ii) are not binary properties; hence there can be a trade-off (e.g., the parameter that allows for better optimization on the current task may be less favourable as a parameter initialization for the next task).
+It is worth noting that such an approach has been successfully applied in the context of online convex optimization in changing environments, where several algorithms can be recognized as running a standard learning algorithm (such as mirror descent) on a carefully selected subset of the parameter space (e.g., Zinkevich, 2003, Herbster and Warmuth, 1998, György and Szepesvári, 2016).
+In this paper, to address problem (i), we first identify key properties that initialization algorithms (Hinton and Salakhutdinov, 2006, He et al., 2015) impose on the starting parameters of a neural network.
+Then, we propose spectral regularization to sustain these properties while striking a good balance between (i) and (ii).
 3.1 Spectral properties at initialization
@@
 3.2 An illustrative example
-To make these points more concrete, consider a single layer neural network $f_{\theta}(x)=\theta_{2}\texttt{ReLU}(\theta_{1}x)$ mapping $x\in\mathbb{R}^{2}$ to $\mathbb{R}$.
+To make these points more concrete, consider a single-layer neural network $f_{\theta}(x)=\theta_{2}\texttt{ReLU}(\theta_{1}x)$ mapping $x\in\mathbb{R}^{2}$ to $\mathbb{R}$.
 Suppose that the first task is to fit $x=(1,0)^{\top}$ and $y=0$ in mean-squared error.
@@
 The condition number of the weight matrix $\theta_{1}$ is $\kappa=1/a$, which requires $\Theta(1/a)=\Theta(\kappa)$ update steps to make significant progress during learning.
-(A more detailed analysis of another example is given in Appendix A.2.)
+(Another example showing this empirically is given in Appendix A.1.)
 3.3 Trainability and Effective Gradient Diversity
@@
 By gradient diversity, we mean the spread of singular vectors in the matrix of per-example stochastic gradients, $\mathbf{G}=[\mathbf{g}_{1},\ldots,\mathbf{g}_{m}]$.
+If this matrix contains a few large singular values and many small singular values, then the gradients will be largely in the span of the singular vectors corresponding to the large singular values.
 For this analysis, we focus on the rank, which is one particular summary statistic for the set of singular values that counts the number of non-zero singular values, $\text{rank}(\mathbf{G})=|\{i:\sigma_{i}(\mathbf{G})>0\}|$.
-However, for practical purposes the rank is problematic because it is unstable to perturbations [34, Theorem 1].
-In our experiments below we will consider the condition number, $\nicefrac{{\sigma_{1}(\mathbf{G})}}{{\sigma_{m}(\mathbf{G})}}$ and effective rank, $\text{erank}(\mathbf{G})=\sum\bar{\sigma}_{i}(\mathbf{G})\log\bar{\sigma}_{i}(\mathbf{G})$ where $\bar{\sigma}_{i}(\mathbf{G})=\nicefrac{{\sigma_{i}(\mathbf{G})}}{{\sum_{i}\sigma_{i}(\mathbf{G})}}$ [35].
-These measures are particularly sensitive to a growing largest singular value, if it grows faster than the smallest singular values (our experiments show this to be the case in Section 5.3).
-In this case, we say there is a reduction in the effective gradient diversity.
-A reduction in the rank of the gradient matrix results in collinear gradients on different datapoints, limiting the diversity of gradient directions used in the updates.
+However, for practical purposes, the rank is problematic because it is unstable to perturbations (Feng et al., 2022, Theorem 1).
+In our experiments below we will consider the condition number, $\nicefrac{{\sigma_{1}(\mathbf{G})}}{{\sigma_{m}(\mathbf{G})}}$ and effective rank, $\text{erank}(\mathbf{G})=\sum\bar{\sigma}_{i}(\mathbf{G})\log\bar{\sigma}_{i}(\mathbf{G})$, where $\bar{\sigma}_{i}(\mathbf{G})=\nicefrac{{\sigma_{i}(\mathbf{G})}}{{\sum_{i}\sigma_{i}(\mathbf{G})}}$ (Roy and Vetterli, 2007).
+These measures make explicit the issues that arise if the largest singular value grows faster than the smallest singular values (our experiments show this to be the case in Section 3).
+In this case, the condition number increases and the effective rank decreases.
+We refer to this as a reduction in the effective gradient diversity.
+A reduction in the rank, or erank, of the gradient matrix results in collinear gradients on different datapoints, limiting the diversity of gradient directions used in the parameter update.
 This can have an adverse effect on learning.
 In the extreme case, where the gradient matrix is rank one, every datapoint provides a gradient in the same direction, even if the datapoints correspond to different classes.
-Consider the gradient of the loss on a particular datapoint $(x_{i},y_{i})$ with respect to the weight matrix of a hidden layer, $\theta_{l}\in\mathbb{R}^{d\times d}$, which can be written recursively as $\mathbf{G}_{l,i}=\nabla_{\theta_{l}}\ell(f_{\theta}(x_{i}),y_{i})=\delta_{l,i}h_{l-1,i}^{\top}$, where $\delta_{l,i}=\theta_{l+1}^{\top}\delta_{l+1,i}\mathbf{D}_{l,i}$ is the error gradient from the previous layer with $\delta_{L,i}=\nicefrac{{\partial\ell(f_{\theta}(x_{i}),y_{i})}}{{\partial f_{\theta}}}$ and $\mathbf{D}_{l,i}=\text{Diag}(\phi^{\prime}([\theta_{l}h_{l,i}]_{1}),\dotso,\phi^{\prime}([\theta_{l}h_{l,i}]_{d}))$.
+Consider the gradient of the loss on a particular datapoint $(x_{i},y_{i})$ with respect to the weight matrix of a hidden layer, $\theta_{l}\in\mathbb{R}^{d\times d}$, which can be written recursively for the parameters of layer $l$ as $\mathbf{G}_{l,i}=\nabla_{\theta_{l}}\ell(f_{\theta}(x_{i}),y_{i})=\delta_{l,i}h_{l-1,i}^{\top}$, where $\delta_{l,i}=\theta_{l+1}^{\top}\delta_{l+1,i}\mathbf{D}_{l,i}$ is the error gradient from the next layer with $\delta_{L,i}=\nicefrac{{\partial\ell(f_{\theta}(x_{i}),y_{i})}}{{\partial f_{\theta}}}$ and $\mathbf{D}_{l,i}=\text{Diag}(\phi^{\prime}([\theta_{l}h_{l,i}]_{1}),\dotso,\phi^{\prime}([\theta_{l}h_{l,i}]_{d}))$.
 We can rewrite the gradient in terms of its vectorization $\mathbf{g}_{l,i}=\text{vec}(\mathbf{G}_{l,i})=(\mathbf{I}_{d}\otimes\theta_{l+1}^{\top})\mathbf{v}_{l,i}$ where $\mathbf{I}_{d}$ is the $d\times d$ identity matrix and ${\mathbf{v}_{1,i}=\text{vec}(\delta_{l+1,i}\mathbf{D}_{l,i}h_{l-1,i}^{\top})\in\mathbb{R}^{d^{2}}}$ is the vectorization of data-dependent terms.
@@
 Thus, if the (effective) rank of $\theta_{l+1}$ decreases, then the (effective) rank of the gradient matrix may decrease as well.
-The rank of the gradient matrix can even decrease due to rank reduction in parameters at other layers, or through the rank of the representation, which others have noted to occasionally correlate with loss of trainability [6, 5, 17].
-3.4 Spectral Properties Away From Initialization During Continual Learning
-It has been observed empirically that the parameter norm grows at a $\sqrt{t}$ rate in large-scale self-supervised learning [36].
-Similar observations have been made in continual learning, provided that the neural network does not stop learning due to loss of trainability [13, 6, 37].
-A growing parameter norm is problematic from an optimisation perspective.
-Specifically, the parameter norm under consideration is the Frobenius norm, equal to the sum of squared singular values, $\|\mathbf{\theta}\|_{F}^{2}=\sum_{i}\sum_{j}[\theta]_{ij}^{2}=\sum_{i}\sigma_{i}(\theta)^{2}$.
-Thus, a growing parameter Frobenius norm is equivalent to an increasing sum of the squared singular values.
-In particular, since $\|\theta_{l}\|_{F}\leq\sqrt{\text{rank}(\theta_{l})}\sigma_{max}(\theta_{l})$ where $\sigma_{max}(\theta_{l})$ denotes the spectral norm (defined as the largest singular value) [38], if the parameter (Frobenius) norm of layer $l$ grows at a $\sqrt{t}$ rate, then the spectral norm of the parameter matrix $\theta_{l}$ also increases at the same rate.
-This leads to an increase in the spectral norm of the layerwise Jacobian, $\sigma_{d}(\mathbf{D}_{l})\sigma_{max}(\theta_{l})\leq\sigma_{max}(\mathbf{J}_{l})$, which may harm trainability.
-Moreover, as we argued in the previous section, an increase in the spectral norm can result in reduced effective gradient diversity.
-4 Regularizing for Trainability in Continual Learning
-If important properties of the initialization are lost during the course of learning, it is natural to regularize the neural network towards the initialization.
-This is the motivation for regularization [17] and weight reinitialization [23, 6] as loss of plasticity mitigators.
+The rank of the gradient matrix can even decrease due to rank reduction in parameters at other layers, or through the rank of the representation, which others have noted to occasionally correlate with loss of trainability (Lyle et al., 2023, Kumar et al., 2023, Dohare et al., 2024).
+Why Do Spectral Properties Deviate From Initialization?
+In large-scale self-supervised learning, it has been observed empirically that the parameter norm grows at a rate of $\sqrt{t}$, where $t$ is the number of iterations (Merrill et al., 2021).
+Similar observations have been made in continual learning, provided that the neural network does not stop learning due to loss of trainability (Nikishin et al., 2022, Lyle et al., 2024, Dohare et al., 2024).
+A growing parameter norm is problematic from an optimization perspective.
+Specifically, the parameter norm that these works consider is the Frobenius norm, which is equal to the sum of squared singular values, $\|\mathbf{\theta}\|_{F}^{2}=\sum_{i}\sum_{j}[\theta]_{ij}^{2}=\sum_{i}\sigma_{i}(\theta)^{2}$.
+Thus, a growing parameter norm is equivalent to an increasing sum of the squared singular values.
+In particular, note that $\|\theta_{l}\|_{F}\leq\sqrt{\text{rank}(\theta_{l})}\sigma_{1}(\theta_{l})$ where $\sigma_{1}(\theta_{l})$ denotes the largest singular value, or the spectral norm (Golub and Van Loan, 2013).
+If the parameter norm of layer $l$ grows at a rate of $\sqrt{t}$, then the spectral norm of the parameter matrix $\theta_{l}$ also increases at the same rate.
+This leads to an increase in the spectral norm of the layerwise Jacobian, $\sigma_{d}(\mathbf{D}_{l})\sigma_{1}(\theta_{l})\leq\sigma_{1}(\mathbf{J}_{l})$, which can reduce effective gradient diversity and may harm trainability.
+4 Spectral Regularization for Continual Learning
+If important properties of the initialization are lost during the course of learning, it is natural to regularize the neural network toward the initialization.
+This is the motivation for regularization (Kumar et al., 2023) and weight reinitialization (Sokar et al., 2023, Dohare et al., 2024) as loss of plasticity mitigators.
 However, our motivation is to more directly target key properties of initialization, using the insights from Section 3.
-To directly preserve key initialization properties, we will develop regularizers $\mathcal{R}_{\tau(t)}(\theta,s)$ that are a function of (i) the parameters of the neural network, $\theta$, (ii) the data, through the current task $\tau(t)$, and of (iii) auxiliary information, such as the parameters at initialization, through the state variable $s$.
+We denote a regularizer by $\mathcal{R}_{\tau(t)}(\theta,s)$, which is a function of (i) the parameters of the neural network, $\theta$, (ii) the data, through the current task $\tau(t)$, and of (iii) auxiliary information, such as the parameters at initialization, through the state variable $s$.
 Naturally, not every regularizer takes all of these elements into account.
@@
 We write the composite objective as $J^{\lambda}_{\tau(t)}(\theta)=J_{\tau(t)}(\theta)+\lambda\mathcal{R}_{\tau(t)}(\theta,s)$, where $\lambda$ is a tunable hyperparameter governing the regularization strength.
+In Section 3, we argued that a growing spectral norm and condition number can harm trainability by reducing effective gradient diversity and that this growth occurs at a rate of $\sqrt{t}$.
+Now we investigate regularization as a means of controlling the spectral norm to maintain the trainability of the neural network.
 One commonly used regularizer in both continual and non-continual learning is L2 regularization, $\mathcal{R}_{\tau(t)}(\theta,\varnothing)=\|\theta\|^{2}$, which regularizes the parameters towards zero.
-A recent alternative, L2 regularization towards the initialization, $\mathcal{R}_{\tau(t)}(\theta,\theta^{(0)})=\|\theta-\theta^{(0)}\|^{2}$, was proposed to deal with sensitivity to parameters near zero [17].
+A recent alternative, L2 regularization towards the initialization, $\mathcal{R}_{\tau(t)}(\theta,\theta^{(0)})=\|\theta-\theta^{(0)}\|^{2}$, was proposed to deal with sensitivity to parameters near zero (Kumar et al., 2023).
 Regularizing towards the particular parameters present at initialization allows the neural network to regenerate parameters, providing a soft reset to parameters if the gradient on the base objective is zero (see Appendix A.2 for an example and Appendix A.4 for more details).
 One potential problem with L2 regularization towards the initialization is that it may prevent the parameters from deviating from the particular sampled value from the initialization distribution.
-4.1 Wasserstein Regularization Towards Initialization
-A first attempt to preserve the properties of the initialization is to regularize the parameter distribution rather than their specific values.
-Thus, we first propose a regularization that is simple and penalizes the parameters if they are far from the initialization distribution.
-At initialization, the parameters at layer $l$ are sampled i.i.d.
-$\theta_{i,j}\sim p^{(l,0)}(\theta)$ according to some pre-determined distribution $p^{(l,0)}$, such as the Glorot initialization [18] The distribution of parameters is only known at the initialization (at iteration 0); for any particular layer, the distribution $p^{(l,t)}$ is no longer has a known closed form at iteration $t$ during training.
-However, it is still possible to regularize the empirical distribution towards the initialization distribution by using the empirical Wasserstein-2 distance, defined as the difference in the order statistics [39].
-We denote the flattened parameter matrix for layer $l$ at iteration $t$ by $\bar{\theta}^{(l,t)}$, and its order statistics by $\theta_{(i)}^{(l,t)}$ for $i=1,\ldots,d_{l}$, representing the $i$-th smallest parameter at iteration $t$ for layer $l$ (here $d_{l}$ denotes the number of parameters at layer $l$).
-The Wasserstein regularizer uses the empirical Wasserstein-2 distance for each layer of the neural network:
-$\mathcal{R}_{\tau(t)}(\theta,\theta^{(0)})=\sum_{l=1}^{L}\sum_{i=1}^{d_{l}}\left(\bar{\theta}_{(i)}^{(l,t)}-\bar{\theta}_{(i)}^{(l,0)}\right)^{2}.$
-In the above equation, we are taking the L2 difference between the order statistics of each layer’s parameters at initialization and at iteration $t$ during training.
-One of the benefits of using this distributional distance is that the parameters have more leeway to move away from the initialization compared to the L2 regularizer (see Appendix A.6 for details).
-4.2 Spectral Regularization for Continual Learning
-In Section 3, we argued that the optimisation dynamics of continual learning may impede the continued trainability of a neural network.
-In particular, we showed that a growing parameter norm results in a growing spectral norm and condition number which can harm trainability by reducing effective gradient diversity.
-Now we investigate regularization as a means of controlling the spectral norm to maintain trainability of the neural network.
-The parameter matrix $\theta_{l}$ in layer $l$ is a composition of the weight matrix $\mathbf{W}_{l}$ and the bias $\mathbf{b}_{l}$, giving the augmented parameter matrix $\theta_{l}=[\mathbf{W}_{l}|\mathbf{b}_{l}]$ (with this the notation the equation defining the of the neural network $f_{\theta}$ can be rewritten as $h_{l+1}(x)=\phi(\mathbf{W}_{l}h_{l}(x)+\mathbf{b}_{l})$).
-As discussed in Section 3, our goal is to control the maximum singular value of $\theta_{l}$.
-Using that $\sigma_{max}(\theta_{l})=\sigma_{max}([\mathbf{W}_{l}|\mathbf{b}_{l}])\leq\sigma_{max}(\mathbf{W}_{l})+\sigma_{max}(\mathbf{b}_{l})$, we achieve this by regularizing the spectral norm of each parameter in the layer separately: the spectral norm of the multiplicative weight parameter $\mathbf{W}_{l}$ is regularized towards 1 and the spectral norm of the additive bias parameter $\mathbf{b}_{l}$ is regularized towards 0:
-$\mathcal{R}_{\tau(t)}(\theta_{l})=\sum_{l\in\text{layers}}\left((\sigma_{max}(\mathbf{W}_{l})^{k}-1)^{2}+\left(\sigma_{max}(\mathbf{b}_{l})^{k}\right)^{2}\right)=\sum_{l\in\text{layers}}\left((\sigma_{max}(\mathbf{W}_{l})^{k}-1)^{2}+\|\mathbf{b}_{l}\|_{2}^{2k}\right).$
+Our proposed spectral regularizer explicitly regularizes each layer’s spectral norm (the maximum singular value) and addresses the reduced effective gradient diversity described in Section 3.3.
+We only minimize the maximum singular value because the smallest singular value remains relatively constant across all the architectures we consider (see Appendix C.6).
+The parameter matrix $\theta_{l}$ in layer $l$ is a concatenation of the weight matrix $\mathbf{W}_{l}$ and the bias $\mathbf{b}_{l}$, giving the augmented parameter matrix $\theta_{l}=[\mathbf{W}_{l}|\mathbf{b}_{l}]$ and augmented input $\tilde{h}_{l}=[h_{l},\mathbf{1}]$.
+With this notation, the equation defining the neural network $f_{\theta}$ can be rewritten as $h_{l+1}(x)=\phi(\mathbf{W}_{l}h_{l}(x)+\mathbf{b}_{l})=\phi(\theta_{l}\tilde{h}_{l})$.
+As discussed in Section 3, our goal is to control the largest singular value of $\theta_{l}$.
+Using the fact that the concatenated parameters can be upper-bounded by the sum, $\sigma_{1}(\theta_{l})=\sigma_{1}([\mathbf{W}_{l}|\mathbf{b}_{l}])\leq\sigma_{1}(\mathbf{W}_{l})+\sigma_{1}(\mathbf{b}_{l})$, we achieve this by regularizing the spectral norm of each parameter in the layer separately: the spectral norm of the multiplicative weight parameter $\mathbf{W}_{l}$ is regularized towards one and the spectral norm of the additive bias parameter $\mathbf{b}_{l}$ is regularized towards zero:
+$\mathcal{R}_{\tau(t)}(\theta_{l})=\sum_{l\in\text{layers}}\left((\sigma_{1}(\mathbf{W}_{l})^{k}-1)^{2}+\left(\sigma_{1}(\mathbf{b}_{l})^{k}\right)^{2}\right)=\sum_{l\in\text{layers}}\left((\sigma_{1}(\mathbf{W}_{l})^{k}-1)^{2}+\|\mathbf{b}_{l}\|_{2}^{2k}\right).$
 We add the exponent $k$ to the spectral norm to penalize large spectral norms deviating from one.
-In our experiments, we set $k=2$ and provide an ablation study in Appendix C.2.
-The maximum singular value can be computed efficiently using power iteration [40, 41].
-Similarly to previous work using spectral regularization for generalization in supervised learning, we find that a single iteration is sufficient for effective regularization [42].
-Spectral regularization explicitly regularizes the spectral norm (the maximum singular value) of each layer and addresses the problem of reduced effective gradient diversity described in Section 3.3.
-We only minimize the maximum singular value.
-We found that the smallest singular value remains relatively constant across all the architectures that we consider (see Appendix C.5).
-For implementation details for other layers, such as normalization layers and convolutional layers, see Appendix A.9.
-In addition to the exponent $k$, our approach to spectral regularization is different in two important ways compared to spectral-norm regularization [42].
+We set $k=2$ in our experiments and provide an ablation study in Appendix C.1.
+For implementation details for other layers, such as normalization and convolutional layers, see Appendix A.7.
+The largest singular value can be computed efficiently using power iteration (Golub and Van der Vorst, 2000, Householder, 2013).
+Similar to previous work using spectral regularization for generalization in supervised learning, we find that a single iteration is sufficient for effective regularization (Yoshida and Miyato, 2017).
+In addition to the exponent $k$, our approach to spectral regularization is different in two important ways compared to previous work, such as spectral-norm regularization (Yoshida and Miyato, 2017).
 First, we regularize the spectral norm of every parameter, including bias terms and normalization parameters (see Appendix A.7).
 This is required because every parameter experiences norm growth.
-Second, we regularize the spectral norm of the multiplicative parameters to one, rather than zero.
+Second, we regularize the spectral norm of the multiplicative parameters to one rather than zero.
 Aggressive regularization of the spectral norm towards zero can lead to collapse issues similar to L2 regularization towards zero.
 These two differences are crucial to improving trainability in continual learning.
-Lastly, spectral regularization is preferable over spectral normalization for two reasons: (i) spectral normalization is data-dependent (see Section 2.1 and Equation 12 in [43], and Appendix A.8), which can be problematic in continual learning due to the changing data distribution, and (ii) other forms of normalization are already used to train deep neural networks, such as LayerNorm [44], and it has been shown that spectral normalization does not improve trainability [5].
+Lastly, spectral regularization is preferable over spectral normalization, which explicitly normalizes the weights in the forward pass to be exactly one, for two reasons: (i) spectral normalization is data-dependent (see Section 2.1 and Equation 12 by Miyato et al.
+(2018), and Appendix A.6), which can be problematic in continual learning due to the changing data distribution, and, (ii) other forms of normalization are already used to train deep neural networks, such as LayerNorm (Ba et al., 2016), and it has been shown that spectral normalization does not improve trainability (Lyle et al., 2023).
 See Appendix A.5 for more details.
 5 Experiments
-Our experiments are designed to investigate the effect of both Wasserstein and spectral regularization on trainability in continual supervised learning.
-They cover several different datasets, types of nonstationarity, and architectures.
-We compare our proposed regularizers against baselines that have shown to improve trainability in previous work, which we detail below.
-Overall, our experiments demonstrate that spectral regularization, and to a lesser extent Wasserstein regularization (i) consistently mitigates loss of trainability on a wide variety of continual supervised learning problems, including training large neural networks for thousands of epochs across hundreds of tasks; (ii) is highly robust to the number of training epochs per task, type of non-stationarity, and choice of regularization strength; and (iii) achieves better generalization performance over the course of continual learning.
+The goal of our experiments is to investigate the effect of spectral regularization on trainability in continual supervised learning, as well as reinforcement learning.
+We cover several different datasets, types of nonstationarity, and architectures.
+We compare our proposed regularizers against baselines that have been shown to improve trainability in previous work, which we detail below.
+Overall, our experiments demonstrate that spectral regularization (i) consistently mitigates loss of trainability on a wide variety of continual supervised learning problems, including training large neural networks for thousands of epochs across a hundred tasks, (ii) is highly robust to the regularization strength, type of non-stationarity and the number of training epochs per task, (iii) achieves better generalization performance over the course of continual learning, and (iv) is generally applicable, which we demonstrate by applying spectral regularization to reinforcement learning with continuous actions.
 Datasets, Nonstationarities, and Architectures
-Our experiments use the following datasets for continual supervised learning: MNIST [45], Fashion MNIST [46], EMNIST [47], SVHN2 [26], CIFAR10, and CIFAR100 [27].
-In addition to the dataset, we consider three different types of non-stationarity: (i) random label assignments, (ii) pixel permutation, and (iii) label flipping.
-Random label assignments are commonly used to evaluate trainability [5] due to the large distribution shift between tasks in memorizing completely new and random labels.
-Pixel permutations, on the other hand, require only learning the permutation mask applied to the image; it induces loss of trainability more slowly but is useful for evaluating generalization [17].
-Lastly, label flipping is a re-assignment of all the observations from one label to another label.
+Our main results uses all commonly used image classification datasets for continual supervised learning: tiny-ImageNet (Le and Yang, 2015), CIFAR10, and CIFAR100 (Krizhevsky, 2009).
+Experiments in the appendix also use smaller-scale datasets, like MNIST (LeCun et al., 1998), Fashion MNIST (Xiao et al., 2017), EMNIST (Cohen et al., 2017), and SVHN2 (Netzer et al., 2011).
+In addition to the dataset, we consider different types of non-stationarity: (i) random label assignments, (ii) pixel permutation, (iii) label flipping, and (iv) class-incremental learning.
+Random label assignments are commonly used to evaluate trainability (Lyle et al., 2023) due to the large distribution shift between tasks in memorizing completely new and random labels.
+Pixel permutations, on the other hand, require only learning the permutation mask applied to the image; it induces loss of trainability more slowly but is useful for evaluating generalization (Kumar et al., 2023).
+Label flipping is a re-assignment of all the observations from one label to another label.
 Unlike the other two non-stationarities, a label flip distribution shift does not require learning a new representation.
-This is because only the output layer of a neural network needs to be permuted to learn the label re-assignment, but gradients can still unnecessarily change the representation leading to loss of trainability [12].
-For the experiments on MNIST, Fashion MNIST and EMNIST, we use feed-forward neural networks.
-On SVHN2, CIFAR10, and CIFAR100 we use both a ResNet-18 [25] and Vision Transformer [24].
+This is because only the output layer of a neural network needs to be permuted to learn the label re-assignment, but gradients can still unnecessarily change the representation leading to loss of trainability (Elsayed and Mahmood, 2024).
+We also consider the class-incremental setting in which the network is trained on a growing subset of the classes from a dataset, starting with only five classes on the first task and introducing five new classes on new tasks.
+We use both a ResNet-18 (He et al., 2016) and Vision Transformer (Dosovitskiy et al., 2021).
 Loss of Trainability Mitigators
-In our main results, we compare Wasserstein regularization and spectral regularization against L2 regularization towards zero, L2 regularization towards the initialization [17], recycling dormant neurons [23], shrink and perturb [15] and no regularization.
-In Appendix C.1, we provide additional results on other baselines: AdamW, which also regularizes towards zero and concatenated ReLU [4, 48].
-Because we are interested in trainability in the task-agnostic setting, we do not compare with regularizers for catastrophic forgetting as they rely on privileged information regarding the task [49, 50].
-Our experiments largely use Layer Norm [44] or Batch Norm in the case of ResNet-18 [51], which has been shown to be a mitigator for loss of trainability [37].
-We also provide an ablation study showing that Layer Norm is synergistic with the regularizers we propose.
+In our main results, we compare spectral regularization against L2 regularization towards zero, shrink and perturb (Ash and Adams, 2020), L2 regularization towards the initialization (Kumar et al., 2023), recycling dormant neurons (ReDO, Sokar et al., 2023), concatenated ReLU (Abbas et al., 2023, Shang et al., 2016), and Wasserstein regularization (Lewandowski et al., 2023).
+Several regularizers in the continual learning without forgetting literature rely on privileged task information, which is not applicable to the task-agnostic setting that we consider.
+We use the streaming conversion (Elsayed and Mahmood, 2024) to transform elastic weight consolidation (Kirkpatrick et al., 2017, Zenke et al., 2017), so that it no longer requires task boundary information, and include it as a baseline.
+Additional experiment details can be found in Appendix B.
 5.1 Comparative Evaluation
-Figure 1: Comparison of task accuracy with random label assignment.
-Spectral regularization is particularly effective at mitigating loss of trainability.
-Without regularization, trainability is always reduced.
-In Figure 1, we plot the results of training a ResNet-18 on different image classification datasets with random label assignment as the type of non-stationarity for each task.
-Across all problems and methods considered, spectral regularization was best at improving and sustaining its trainability, with accuracy increasing as new tasks were encountered.
-In contrast, despite improved trainability over the first few tasks, the unregularized baseline eventually experienced loss of trainability, leading to lower accuracy.
-Recycling dormant neurons, which reinitializes weights based on dormant neurons, experienced loss of trainability on all problems.
-One potential reason for this is because this approach only resets the weights of the convolutional layers and the final fully-connected layer [23], whereas our approach regularizes all parameters, including bias terms and batch normalization parameters.
-In Appendix C.1, we show that weight reinitialization can be effective on smaller problems without normalization layers.
-Shrink and perturb is similar to L2 regularization towards zero, but injects noise and is not as effective.
-While the perturbations were previously demonstrated to improve generalization when warm-starting a neural network on a new problem, we see that they are not effective when training on a longer sequence of problems.
-Wasserstein regularization is often on par with L2 regularization towards the initialization, and surpasses it on CIFAR10.
-These results demonstrate that, in contrast to other approaches, spectral regularization consistently maintains trainability across problems.
+Figure 1: Generalization across different types of non-stationarity on tiny-ImageNet using a ResNet (top) or a Vision Transformer (bottom).
+Compared to the baselines, spectral regularization is consistently among the best-performing methods across class incremental, label flip, and pixel permutation non-stationarities.
+Note that the Vision Transformer often achieves better generalization performance than the ResNet architecture.
+In Figure 1, we plot the results of training a ResNet-18 and a Vision Transformer on tiny-ImageNet with different non-stationarities.
+Across networks, non-stationarities, and methods considered, we see that spectral regularization is among the methods best capable of sustaining plasticity.
+The advantage of our approach, spectral regularization, is particularly high in class-incremental learning, but the performance of spectral regularization with a Vision Transformer was also the best on label flipping and pixel permutation.
+In contrast, the performance of other baselines was highly variable with respect to the specific type of non-stationarity.
+For example, shrink and perturb is not robust to the settings considered.
+Sometimes, it is among the best-performing or worst-performing methods.
+Similar results on other datasets and the pixel permutation non-stationarity can be seen in Figure 2.
+Figure 2: Continual learning with pixel permutation tasks on SVHN2, CIFAR10, CIFAR100 using a ResNet-18 (top) or a Vision Transformer (bottom).
+Across different datasets, spectral regularization is effective at maintaining test accuracy on new tasks.
+Without any mitigators, both ResNet-18 and Vision Transformer have diminishing test accuracy, suggesting loss of plasticity.
+5.2 Looking Inside the Network
+Figure 3: Trainability and neural network properties across ImageNet, CIFAR10, and CIFAR100.
+Baselines that suffer from a loss of trainability (top) also have an increasing average spectral norm (middle-top), and a decrease in their average representation change (middle-bottom).
+We now explore how the structural properties of a neural network evolve over the course of continual learning, how these properties are affected by spectral regularization, and the different baselines considered.
+For this, we consider the continual learning problem in which a ResNet-18 must memorize a set of random labels which changes from task to task using the tiny-ImageNet, CIFAR10, and CIFAR100 datasets.
+We consider the average representation change to measure the distance between the neural network’s hidden activations from the beginning of one task, to the beginning of the next task.
+The average representation change is a proxy for plasticity, allowing us to see how much the behavior of the neural network has changed.
+In Figure 3, we see that the unregularized ReLU networks suffers from loss of trainability in all problems considered, and this coincides with an increasing spectral norm (middle-top) and a decrease in the average representation change (bottom).
+Although only spectral regularization directly regularizes the spectral norm of the network, other regularizers do so indirectly by controlling other norms, like L2.
+However, these other regularizers also regularize other parameters, preventing them from deviating from initialization and potentially leading to suboptimal use of capacity, which can be observed in the bottom row in Figure 3.
+In Appendix C.3, we show that the Vision Transformer was not able to memorize random labels, suggesting that its capabilities for generalization are offset by its relatively lower trainabilty.
 5.3 Sensitivity Analysis
-In this section, we investigate the robustness of spectral regularization when varying (i) the number of training epochs per task, (ii) the type of non-stationarity applied, or, (iii) the regularization strength.
-Varying the Number of Epochs Per Task
-Given that the continual learning problem outlined in Section 2 depends on the number of iterations, it may be the case that a higher number of epochs per task can mitigate loss of trainability.
-In Figure 2, we found that the number of epochs only delays loss of trainability (see Appendix C.3 for results on Fashion MNIST).
-Even when the number of epochs per task is high enough to reach 100% accuracy on the first task, loss of trainability eventually occurred without regularization.
-In contrast, when using spectral regularization, loss of trainabililty was consistently mitigated.
-We additionally found that spectral regularization is particularly effective at maintaining trainability when the number of epochs per tasks is low.
-The evidence for this is most striking without Layer Norm, in the top row, where spectral regularization was the only method capable of maintaining its trainability.
-Even though Layer Norm controls the spectral norm through the activations [52], adding Layer Norm alone was not enough to mitigate loss of trainability.
-However, Layer Norm is synergistic with various forms of regularization, with accuracy being improved regardless of the number of epochs per task.
-Figure 2: Sensitivity to number of epochs per task on MNIST with random label assignment.
-Without Layer Norm, spectral regularization is able to maintain trainability over other regularizers even when the number of epochs is low.
-Spectral regularization also synergizes with Layer Norm, improving initial performance on the first task, and improving over tasks.
-Figure 3: Sensitivity to type of non-stationarity on EMNIST.
-Both Wasserstein and spectral regularization are able to consistently maintain high trainability across random label assignment, pixel permutations, and label flipping.
-When applying random label assignment, L2 (init) and L2 (zero) are unable to attain high trainability as compared to Wasserstein and spectral regularization.
-Varying Type of Non-stationarity
-Next, we explore whether our proposed regularizers are effective across different types of non-stationarity.
-To this end, we use EMNIST because the number of classes is large enough that label flipping induces loss of trainability [12].
-In Figure 3, we found that both Wasserstein and spectral regularization best maintained trainability across different non-stationarities.
-Figure 4: Sensitivity to regularization strength on MNIST with random label assignment. Spectral regularization is insensitive to regularization strength while remaining performant at few epochs.
-Varying Regularization Strength
+Figure 4: Sensitivity analysis on regularization strength.
+Compared to other regularizers, spectral regularization is insensitive to regularization strength while sustaining higher trainability for any given regularization strength.
 The effectiveness of a regularizer depends on the regularization strength.
@@
 However, too little regularization can lead to loss of trainability and suboptimal performance on future tasks.
-In a continual learning problem, hyperparameter tuning is particularly costly (or infeasible), especially if the distribution over tasks is not known in advance [53].
+In a continual learning problem, hyperparameter tuning is particularly costly (or infeasible), especially if the distribution over tasks is not known in advance (Mesbahi et al., 2024).
 Thus, we should favour regularizers that are less sensitive to hyperparameters.
-Our results in Figure 4 show that spectral regularization dominates other regularizers for any choice of regularization strength, is much highly insensitive to the choice of hyperparameter compared to other regularizers.
-Figure 5: Effects of regularization on distance from initialization, average condition, average weight rank, and gradient diversity when training ResNet-18 on SVHN2 with random label assignment.
-Despite being far from initialization, spectral regularization has a low average condition number, a high effective rank across its parameters, and a high effective gradient diversity.
-5.3 Looking Inside the Network
-We now explore structural properties of the neural network, and how these properties are effected by our proposed regularizers.
-In Figure 5, we further confirm that spectral regularization is the least restrictive approach, allowing deviations far from the initialization.
-Spectral regularization also controlled the spectral norm of the network, leading to a better average condition number (the ratio between the maximum and the minimum singular values), a higher effective rank of its parameters, and a higher gradient diversity.
-See Appendix C.5 for similar results on the other datasets considered.
-Figure 6: Comparison of generalization performance with random label assignment. Spectral regularization improves generalization, using either a ResNet-18 (Top) or a Vision Transformer (bottom).
-5.4 From Trainability to Generalization
-Having demonstrated the main goal of this paper, namely the effectiveness of Wasserstein and spectral regularization on trainability, we now investigate generalization.
-We used a pixel permutation non-stationarity in Figure 6, and found that spectral regularization is better able to generalize compared to other regularizers, using either a ResNet-18 or Vision Transfomer.
-Interestingly, we also find that the Vision Transformer, particularly with spectral regularization, provides better generalization performance than using a ResNet.
-This demonstrates that regularizers which more directly address the issues underlying trainability also lead to better generalization.
+Our results in Figure 10 show that the sensitivity of spectral regularization to its hyperparameter is much lower than other regularizers.
+In Appendix C.4, we also present additional results showing the robustness of spectral regularization when varying (i) the regularization strength using other network architectures, (ii) the type of non-stationarity, and (iii) the number of training epochs per task.
+Refer to captionRefer to caption
+Refer to captionRefer to caption
+Refer to captionRefer to captionRefer to caption
+Figure 5: Spectral regularization enhances plasticity in reinforcement learning in the DMC suite.
+Spectral regularization is competitive with the network reset + layernorm (Reset) even when the replay buffer is unbounded (Top Left).
+When the replay buffer size is bounded to 250k steps, spectral regularization improves over the Reset baseline (Bottom Left).
+In both cases, Spectral regularization significantly outperforms layernorm (Baseline), Compared to the hyperparameter governing the reset frequency, spectral regularization is less sensitive to its regularization strength.
+Spectral regularization also prevents both parameter and gradients from exploding, and reduces value overestimation (Right).
+5.4 From Supervised Learning to Reinforcement Learning
+In addition to supervised learning, we evaluate spectral regularization in reinforcement learning.
+Specifically, we investigate the control tasks from the DMC benchmark (Tassa et al., 2020), with soft actor-critic  (SAC, Haarnoja et al., 2018), in the high replay-ratio (RR) regime (D’Oro et al., 2022) with 16 gradient updates for every environment step.
+The high replay-ratio regime leads to primacy bias (Nikishin et al., 2022), a phenomenon related to loss of plasticity.
+Recent work has demonstrated that only full network resets with layer normalization (Ba et al., 2016) serve as an effective mitigation strategy in this setting (Ball et al., 2023, Nauman et al., 2024).
+As shown in Figure 5, spectral regularization yields substantially higher returns compared to the baseline that uses layer normalization with either an unbounded replay buffer size or a limited replay buffer size.
+In aggregate, spectral regularization is competitive with the final performance of the strongest baseline, full network resets in combination with layer normalization.
+However, in five out of seven environments, spectral regularization substantially improves over the reset baseline in terms of sample efficiency (see Appendix C.8).
+Spectral regularization is also less sensitive to its regularization strength than resets are for their reset frequency (Figure 5, middle-top and middle-bottom).
+Moreover, we used a single regularization strength for every network.
+Better performance may be achieved by individually tuning each regularization strength for the value and policy networks.
+In addition, a combination of spectral normalization and layer norm aggressively reduces Q-value overestimation (Figure 5, top-left), which is one of the well-established proxies for RL training destabilization (Hasselt, 2010, Nauman et al., 2024).
+Lastly, spectral normalization prevents exploding gradients and keeps weights’ norms small during training, which is a key component for learning continually and preserving the plasticity of neural networks (Lyle et al., 2023; 2022).
 6 Conclusion
-In this paper, we investigated the connection between initialization, trainability and regularization.
-We identified that controlling the singular values so that they do not deviate significantly from their values at initialization promotes diverse gradients for neural networks that sustains trainability throughout continual learning.
-As the first approach to preserve trainability properties induced by initialization, building on recent work, we proposed to regularize the distribution of individual parameters during training towards the distribution at initialization.
-We then proposed to use spectral regularization as a way of controlling the spectral norm of each layer, thereby sustaining the trainability properties that are present at initialization.
-Our experiments show that spectral regularization is more performant and less sensitive to hyperparameters than other regularizers across datasets, nonstationarities, and architectures.
+In this paper, we investigated the connection between initialization, trainability, and regularization.
+We identified that deviations of the maximum singular value of each layer can lead to low gradient diversity, preventing neural networks from training on new tasks.
+To directly preserve trainability properties present at initialization, we proposed spectral regularization as a way of controlling the maximum singular value of each layer so that it does not deviate significantly from $1$.
+Our experiments show that spectral regularization is more performant and less sensitive to hyperparameters than other methods across datasets, nonstationarities, and architectures.
 While several methods mitigate loss of trainability to a varying degree, learning continually with spectral regularization is robust, achieving high initial and sustained performance.
-Finally, we show that spectral regularization is also capable of improving generalization with both Vision Transformer and ResNet-18 on continual versions of SVHN2, CIFAR10 and CIFAR100.
+We also showed that spectral regularization is capable of improving generalization with both Vision Transformer and ResNet-18 on continual versions of tiny-ImageNet, CIFAR10, CIFAR100, amongst others.
+Lastly, we showed that spectral regularization learned a more effective policy in reinforcement learning by avoiding early loss of plasticity due to primacy bias.
 Appendix A Additional Details
-A.1 Common Assumptions Regarding Tasks
-We denote the parameters of a randomly initialized neural network trained only on task $\tau$ for $T$ iterations as $\theta_{\tau}^{(T)}=\theta^{(T-1)}_{\tau}-\alpha\nabla_{\theta}J_{\tau}(\theta)\big{|}_{\theta=\theta_{\tau}^{(t-1)}}$, with $\theta^{(0)}_{\tau}$ being sampled from a standard initialization distribution, such as Glorot [18] or He initialization [29].
-The equal-difficulty assumption can be formalized by assuming that the objective value, $J_{\tau}(\theta_{\tau}^{(T)})$, is statistically independent of $\tau$.
-That is, while the data depends on the task, the objective value after training does not depend on the task.
-Note that the equal-difficulty assumption is key: otherwise it would be difficult to disambiguate whether a performance decrease is due to a more difficult task or due to loss of trainability in the learning algorithm.
-A.2 Illustrative Example: Large Spectral Norm Can Impede Trainability
+A.1 Illustrative Example: Large Spectral Norm Can Impede Trainability
 We show that, given a family of solutions on the first task, preference should be given to solutions that have a spectral norm that is closer to one.
@@
 We consider the squared loss, and use a mutli-layer perceptron with one hidden layer and two units, $f_{w_{2},w_{1}}(x)=w_{2}\texttt{ReLU}(w_{1}x)$.
-A common observation regarding solutions found for deep neural networks is that they are low rank [54, 55, 56, 57].
-In addition, it has been observed that large outliers in the spectrum of a deep network emerge during training, meaning that the largest singular values of the weight matrices tend to be much larger than their smallest singular values [58, 22].
+A common observation regarding solutions found for deep neural networks is that they are low rank (Arora et al., 2019, Jacot, 2023, Timor et al., 2023, Galanti et al., 2022).
+In addition, it has been observed that large outliers in the spectrum of a deep network emerge during training, meaning that the largest singular values of the weight matrices tend to be much larger than their smallest singular values (Pennington et al., 2018, Martin and Mahoney, 2021).
 Consider one such solution where $w_{1}=\begin{bmatrix}-1&c\\ \nicefrac{{1}}{{c}}&-1\end{bmatrix}$, and $w_{2}=[\nicefrac{{1}}{{c}},c]\in\mathbb{R}^{1\times 2}$.
@@
 where $\ell_{1}=\ell(f_{w_{2},w_{1}}(x_{1}),y^{\prime}_{1})$, and $\alpha$ is a step-size.
-One of the consequences of the imbalanced singular values is that the per-example gradients are poorly conditioned [59].
+One of the consequences of the imbalanced singular values is that the per-example gradients are poorly conditioned (Wu et al., 2021).
 That is, when we sample $x_{1}$, there will be a large gradient for $w_{2}$, $\|\nabla_{w_{2}}\ell(f_{w_{2},w_{1}}(x_{1}),y_{1}^{\prime})\|^{2}=c$ but a small gradient for $w_{1}$, $\|\nabla_{w_{1}}\ell(f_{w_{2},w_{1}}(x_{1}),y_{1}^{\prime})\|^{2}=\nicefrac{{1}}{{c}}$.
@@
 Figure 6: Illustrative Example. A lower spectral norm is more trainable across different step sizes in the simple illustrative example.
-A.3 Differences Between Continual and Non-continual Learning
+A.2 Differences Between Continual and Non-continual Learning
 One critical difference between continual and non-continual learning is that, in continual learning, convergence towards a fixed point is not the objective.
@@
 The neural network will reach a fixed point in continual learning, meaning that the gradients are zero for every datapoint, if either 1) both $\theta_{1}$ and $\theta_{2}$ are zero, or 2) the activations are zero for every input $x$, $\phi(\theta_{1}x)=0$,
-A.4 Spectral Norm and Lipschitz Constant
+A.3 Spectral Norm and Lipschitz Constant
 One other potential reason for why a growing spectral norm is problematic is from the theory convex optimisation.
 For the ReLU activation function, the Lipschitz constant of the layer map, $h_{l+1}=\phi(\theta_{l}h_{l})$, is equal to the spectral norm of the weight matrix.
-We can thus bound the Lipschitz constant of the entire neural network above by the product of the layer-wise Lipschitz constants [60].
+We can thus bound the Lipschitz constant of the entire neural network above by the product of the layer-wise Lipschitz constants (Szegedy et al., 2013).
 Putting these results together, we conclude that during continual learning, the Lipschitz constant of each layer is increasing and thus the Lipschitz constant of the entire network is also increasing.
@@
 Thus, if the Lipschitz constant grows then, eventually, the chosen step-size will be too large and gradient descent will not converge locally on the task.
-A.5 More Details on Regularization and Trainability
-One other important advantage of regularization is that it has well understood effects for ensuring trainability in approaches outside of deep learning, such as degenerate linear regression and inverse problems [61].
+A.4 More Details on Regularization and Trainability
+One other important advantage of regularization is that it has well understood effects for ensuring trainability in approaches outside of deep learning, such as degenerate linear regression and inverse problems (Benning and Burger, 2018).
 Each regularizer biases the parameter dynamics in their own way, such as keeping the parameters close to zero or close to initialization.
@@
 But, with regularization, the weights will be updated and the unit may become active again for some inputs.
-A.6 More Details on Wasserstein Regularizer
-At initialization, the parameters at layer $l$ are sampled i.i.d.
-$\mathbf{\theta}_{i,j}\sim p^{(l,0)}(\theta)$ according to some pre-determined distribution, such as the Glorot initialization [18].
-The distribution of parameters at iteration $t$ during training and for any particular layer, denoted by $p^{(l,t)}$, is no longer known and does not have a closed form.
-However, it is still possible to regularize the empirical distribution towards the initialization distribution by using the empirical Wasserstein metric [39].
-We denote the flattened parameter matrix for layer $l$ at time $t$ by $\mathbf{\bar{\theta}}^{(l,t)}$.
-The squared Wasserstein-2 distance between the distribution of parameters at initialization and the current parameter distribution is defined as
-$\mathcal{R}_{\tau(t)}(\theta,\theta^{(0)})=\sum_{i=1}^{d}\left(\bar{\theta}_{(i)}^{(l,t)}-\bar{\theta}_{(i)}^{(l,0)}\right)^{2}.$
-Unlike the regenerative regularizer, the Wasserstein regularizer takes the difference of the order statistics.
-Thus, the regenerative regularizer is always larger because the Wasserstein regularizer takes the difference in the sorted values, $\sum_{i=1}^{d}\left(\mathbf{\bar{\theta}}_{(i)}^{(l,t)}-\mathbf{\bar{\theta}}_{(i)}^{(l,0)}\right)^{2}\leq\sum_{i=1}^{d}\left(\mathbf{\bar{\theta}}_{i}^{(l,t)}-\mathbf{\bar{\theta}}_{i}^{(l,0)}\right)^{2}$.
-The Wasserstein regularizer, like variants of L2, regularize the individual parameters and may not preserve the properties of the entire parameter matrix at each layer.
-A.7 Details Regarding Spectral Regularization
+A.5 Details Regarding Spectral Regularization
 Spectral Regularization vs L2 Regularization
-L2 regularization is not as effective as spectral regularization because it constrains the magnitude of the parameters, and can even cause rank collapse [17].
+L2 regularization is not as effective as spectral regularization because it constrains the magnitude of the parameters, and can even cause rank collapse (Kumar et al., 2023).
 Spectral regularization constrains only the spectral norm of the weight matrix close to 1, $\sigma^{(1)}(\theta_{l})\approx 1$.
@@
 Spectral Regularization vs Spectral Normalization:
-The spectral norm can also be controlled via spectral normalization [43].
-For continual learning, spectral regularization is preferable over spectral normalization for two reasons: (i) spectral normalization is data-dependent (see Section 2.1 and Equation 12 in [43], and Appendix A.8 for more details), which can be problematic in continual learning due to the changing data distribution, and, (ii) other forms of normalization are often already used to train deep neural networks, such as LayerNorm [44].
+The spectral norm can also be controlled via spectral normalization (Miyato et al., 2018).
+For continual learning, spectral regularization is preferable over spectral normalization for two reasons: (i) spectral normalization is data-dependent (see Section 2.1 and Equation 12 in Miyato et al.
+(2018), and Appendix A.6 for more details), which can be problematic in continual learning due to the changing data distribution, and, (ii) other forms of normalization are often already used to train deep neural networks, such as LayerNorm (Ba et al., 2016).
 LayerNorm, in particular, regularizes the magnitude of the layerwise map by dividing by the norm of the features in the layer.
@@
 That being said, normalization can still improve performance and optimal performance may be achieved by a combination of regularization and normalization, which we show in Section 5.3.
-A.8 Categorizing Regularizers for Continual Learning
+A.6 Categorizing Regularizers for Continual Learning
 We can categorize regularizers broadly into data-dependent and data-indepedent regularizers.
@@
 Regularizing on one distribution of data does not necessarily maintain useful properties for learning on the next distribution of data.
-Examples of data-dependent regularizers include feature rank regularization [62], auxiliary tasks and other feature-space regularizers [14].
-Other non-standard approaches that have a data-dependent regularization effect include gradient regularizers such as gradient-clipping and weight reinitialization based on dormant neurons [23].
+Examples of data-dependent regularizers include feature rank regularization (Kumar et al., 2021), auxiliary tasks and other feature-space regularizers (Lyle et al., 2022).
+Other non-standard approaches that have a data-dependent regularization effect include gradient regularizers such as gradient-clipping and weight reinitialization based on dormant neurons (Sokar et al., 2023).
 These data-dependent regularizers regularizers control some property on data from the current task but may not control the property on data from a new task, which is needed to maintain plasticity.
@@
 This category regularizes the parameters directly, which is particularly useful in continual learning when the data distribution is changing.
-Examples of data-independent regularization include L2 regularization towards zero, weight decay (for adaptive gradient methods), and regenerative regularization [17].
+Examples of data-independent regularization include L2 regularization towards zero, weight decay (for adaptive gradient methods), and regenerative regularization (Kumar et al., 2023).
 Our proposed approach of spectral regularization is data-independent and, moreover, explicitly targets a trainability condition for continual learning.
 Thus, we expect that it is particularly effective on maintaining trainability in continual learning compared to other data-independent regularizers.
-A.9 Spectral Regularization of Other Layers
+A.7 Spectral Regularization of Other Layers
 Normalization Layers
-Other layers using per-unit scaling, such as normalization layers [51, 44], are also spectrally regularized.
-For example, denote $\gamma\in\mathbb{R}^{d}$ as the trainable scaling parameters for LayerNorm [44], then the element-wise product of the weights can be written as a diagonal weight matrix, $f(z)=\gamma\circ z=\text{Diag}(\gamma)z$.
+Other layers using per-unit scaling, such as normalization layers (Ioffe and Szegedy, 2015, Ba et al., 2016), are also spectrally regularized.
+For example, denote $\gamma\in\mathbb{R}^{d}$ as the trainable scaling parameters for LayerNorm (Ba et al., 2016), then the element-wise product of the weights can be written as a diagonal weight matrix, $f(z)=\gamma\circ z=\text{Diag}(\gamma)z$.
 The maximum singular value of a diagonal matrix is the maximum entry, $\sigma_{max}(\text{Diag}(\gamma))=\max_{i}|\gamma_{i}|$.
@@
 Convolutional Layers
-Similar to other work, we reshape the convolutional weight tensor (with kernel size ($k\times k$, $d_{in}$ filters and $d_{out}$ filters) into a matrix of size $d_{out}\times(k\cdot k\cdot d_{in})$ [42].
-The spectral norm of this reshaped matrix provides an efficient upper bound on the spectral norm of the Toeplitz matrix defining the convolution [63, Corollary 1].
+Similar to other work, we reshape the convolutional weight tensor (with kernel size ($k\times k$, $d_{in}$ filters and $d_{out}$ filters) into a matrix of size $d_{out}\times(k\cdot k\cdot d_{in})$ (Yoshida and Miyato, 2017).
+The spectral norm of this reshaped matrix provides an efficient upper bound on the spectral norm of the Toeplitz matrix defining the convolution (Tsuzuku et al., 2018, Corollary 1).
 Appendix B Experiment Details
-All of our experiments used Adam [64] where the default step size of $0.001$ was selected after an initial sweep over $[0.005,0.001,0.0005]$.
+All of our experiments used Adam (Kingma and Ba, 2015) where the default step size of $0.001$ was selected after an initial sweep over $[0.005,0.001,0.0005]$.
 For all of our results, we use 10 random seeds and provided a shaded region corresponding to the standard error of the mean.
-For experiments on SVHN2, CIFAR10 and CIFAR100, we used 5 seeds to sweep over the regularization strengths of $[0.01,0.001,0.0001]$, and found that $0.0001$ worked well on CIFAR10 and CIFAR100 for all regularizers, whereas $0.001$ worked best on SVHN2 for all regularizers.
+For experiments on tiny-ImageNet SVHN2, CIFAR10 and CIFAR100, we used 4 seeds to sweep over the regularization strengths of $[0.01,0.001,0.0001]$, and found that $0.0001$ worked well on tiny-ImageNet, CIFAR10 and CIFAR100 for all regularizers, whereas $0.001$ worked best on SVHN2 for all regularizers.
 Datasets and non-stationarities:
@@
 We used a batch size of 250 and found this to be effective.
-For random label non-stationarity, 20 epochs per task and 25 tasks was enough to lose trainability for an unregularized network.
+For random label non-stationarity, 20 epochs per task and 30 tasks was enough to lose trainability for an unregularized network.
 For pixel permutation non-stationarity, 10 epochs and 100 tasks was enough to lose trainability for an unregularized network.
@@
 We used a batch size of 250 and found this to be effective.
-For random label non-stationarity, 20 epochs per task and 25 tasks was enough to lose trainability for an unregularized network.
+For random label non-stationarity, 20 epochs per task and 30 tasks was enough to lose trainability for an unregularized network.
 For pixel permutation non-stationarity, 10 epochs and 100 tasks was enough to lose trainability for an unregularized network.
+  • tiny-ImageNet: All of the 100000 images were used for training, 10000 images were used for validation, and 10000 images were used for testing according to the predetermined split.
+We used a batch size of 250 and found this to be effective.
+For random label non-stationarity, 20 epochs per task and 30 tasks was enough to lose trainability for an unregularized network.
+For pixel permutation non-stationarity, 20 epochs and 100 tasks was enough to lose trainability for an unregularized network.
+  • Reinforcement Learning: We evaluate spectral regularization in RL control tasks from DMC benchmark (Tassa et al., 2020), with SAC method (Haarnoja et al., 2018) in demanding replay ratio (RR) regime (D’Oro et al., 2022) with 16 gradient updates per every new environments step.
+We choose this setup because a high RR regime leads to significant primacy bias (Nikishin et al., 2022) defined as a tendency to overfit initial experiences that damages the rest of the learning process.
 Neural Network Architectures:
-  • For MNIST, EMNIST, and Fashion MNIST: 4-layer MLP with 256 neurons per layer and relu activations.
+  • MNIST, EMNIST, and Fashion MNIST: 4-layer MLP with 256 neurons per layer and relu activations.
 For experiments that use it, LayerNorm was applied after the linear weight matrix and before the non-linearity.
-  • For SVHN2, CIFAR10, and CIFAR100: An off-the-shelf ResNet-18 with batch norm, as well as an off-the-shelf Vision Transformer (tiny).
+  • tiny-ImageNet CIFAR10, CIFAR100, and SVHN2: An off-the-shelf ResNet-18 with batch norm, as well as an off-the-shelf Vision Transformer (tiny).
+  • Reinforcement Learning: Recent studies (Nauman et al., 2024, Ball et al., 2023) have demonstrated that, in this setup, only resets with layer normalization (Ba et al., 2016) serve as an effective mitigation strategy.
+We compare the performance of the SAC agent with both spectral regularization and layer normalization to two baseline agents: SAC with only layer normalization, and SAC with layer normalization plus resets.
+We use spectral regularization with a coefficient $41e-$ 4 for both the actor and critic.
+For every method, we use a single critic and architecture size of $2$ layers and $256$ neurons per layer for both actor and critic.
+Using a random policy, we prefill a replay buffer with $10,000$ transitions before starting the training.
+Replay buffer maximum size is $1$ million transitions.
 Appendix C Additional Experiments
-C.1 Other Baselines
-Figure 8: AdamW in Random Label MNIST: AdamW performs similarly to L2 regularization towards zero, recovering trainability when LayerNorm is used.
-Figure 9: Concatenated ReLU, Random Label MNIST. CReLU is unable to maintain trainability even with layer normalization.
-Figure 10: Recycling Dormant Neuron, Random Label MNIST.
-Recycling dormant neurons are effective without layer normalization, but layer normalization removes dormant neurons and still leads to loss of trainability.
-C.2 Ablating Hyperparameter $k$ in Spectral Regularization
-Figure 11: Evaluating the choice of $k$ for spectral regularization for $k=1,2,4,8$ (Left) MNIST, (Right) Fashion MNIST.
-C.3 Results on Fashion MNIST
-Figure 12: Loss of Plasticity in Fashion MNIST Regularization is necessary, LayerNorm without regularization loses plasticity: the performance difference increases as a function of tasks, even with an increasing number of iterations.
-C.4 Effect of Regularization on Capacity in a Single Task
+C.1 Ablating Hyperparameter $k$ for Spectral Regularization
+Figure 7: Evaluating the choice of $k$ for spectral regularization for $k=1,2,4,8$ on MNIST (Right) and Fashion MNIST (left).
+We found that $k=2$ balance stability with effectiveness, and use this value throughout our experiments.
+C.2 Results on Fashion MNIST
+Figure 8: Loss of Plasticity in Fashion MNIST.
+Although the all the newtorks in this experiment use LayerNorm, loss of plasticity occurs without regularization: the performance decreases as a function of tasks, even with an increasing number of iterations.
+C.3 Vision Transformer Cannot Memorize Random Labels
+Figure 9: ResNet and Vision Transformer on the random label memorization task using tiny-ImageNet.
+The Vision Transformer architecture is particularly well-suited to tasks with structure from which generalization is possible.
+However, we found that its trainability on random labels is lacking.
+We were unable to get the Vision Transformer to memorize random labels even on a single task in ImageNet.
+C.4 Additional Sensitivity Results
+Figure 10: Sensitivity analysis on regularization strength Compared to other regularizers, spectral regularization is insensitive to regularization strength while sustaining higher trainability for any given regularization strength.
+Robustness to Non-stationarity in Fashion MNIST
+Next, we explore whether our proposed regularizers are effective across different types of non-stationarity.
+To this end, we use EMNIST because the number of classes is large enough that label flipping induces loss of trainability (Elsayed and Mahmood, 2024).
+In Figure 11, we found that spectral regularization best maintained trainability across different non-stationarities.
+Figure 11: Sensitivity to type of non-stationarity on EMNIST.
+Spectral regularization is able to consistently maintain high trainability across random label assignment, pixel permutations, and label flipping.
+When applying random label assignment, L2 (init) and L2 (zero) are unable to attain high trainability as compared to spectral regularization.
+Varying the Number of Epochs Per Task
+Given that the continual learning problem outlined in Section 2 depends on the number of iterations, it may be the case that a higher number of epochs per task can mitigate loss of trainability.
+In Figure 12, we found that the number of epochs only delays loss of trainability (see Appendix C.2 for results on Fashion MNIST).
+Even when the number of epochs per task is high enough to reach 100% accuracy on the first task, loss of trainability eventually occurred without regularization.
+In contrast, when using spectral regularization, loss of trainabililty was consistently mitigated.
+We additionally found that spectral regularization is particularly effective at maintaining trainability when the number of epochs per tasks is low.
+The evidence for this is most striking without Layer Norm, in the top row, where spectral regularization was the only method capable of maintaining its trainability.
+Even though Layer Norm controls the spectral norm through the activations (Kim et al., 2021), adding Layer Norm alone was not enough to mitigate loss of trainability.
+However, Layer Norm is synergistic with various forms of regularization, with accuracy being improved regardless of the number of epochs per task.
+Figure 12: Sensitivity to number of epochs per task on MNIST with random label assignment.
+Without Layer Norm, spectral regularization is able to maintain trainability over other regularizers even when the number of epochs is low.
+Spectral regularization also synergizes with Layer Norm, improving initial performance on the first task, and improving over tasks.
+C.5 Effect of Regularization on Capacity in a Single Task
 The regularization strength for preventing loss of trainability must be sufficiently high, and this regularization strength may limit the capacity of the neural network.
@@
 The unregularized baseline is able to use its capacity fully, but at the cost of reduced trainability in later tasks.
-C.5 Investigating Effects of Regularization on Neural Network Properties
-Figure 14: Singular Values and Weight Norms: Top: SVHN2, Middle: CIFAR10, Bottom: CIFAR100.
-Figure 15: Rank and Singular Value Spread: Top: SVHN2, Middle: CIFAR10, Bottom: CIFAR100.
-C.6 Investigating Generalization on Continual Imagenet
+C.6 Investigating Effects of Regularization on Neural Network Properties
+Figure 14: Singular values, weight norms, stable ranks and effective gradient diversity: Left: SVHN2, Middle: CIFAR10, Right: CIFAR100.
+C.7 Investigating Generalization on Continual Imagenet
 We investigate the generalization performance of spectral regularization on Continual ImageNet.
 In Continual ImageNet, each task is to distinguish between two ImageNet classes.
-We use the network architecture and training protocol used in [17].
+We use the network architecture and training protocol used in Kumar et al.
+(2023).
 We find that all regularization approaches tested achieve high generalization performance across tasks.
@@
 Figure 15: Generalization performance on Continual ImageNet.
+C.8 Results on Individual DMC Environments
+We report mean returns for DMC environments in Figure 16.
+Figure 16: Mean and standard error of return for 7 DMC environments.